### v1

# Codimension-Two Spiral Spin-Liquid in the Effective Honeycomb-Lattice Compound Cs3Fe2Cl9

Abstract
A codimension-two spiral spin-liquid is a correlated paramagnetic state with one-dimensional ground state degeneracy hosted within a three-dimensional lattice. Here, via neutron scattering experiments and numerical simulations, we establish the existence of a codimension-two spiral spin-liquid in the effective honeycomb-lattice compound Cs3Fe2Cl9 and demonstrate the selective visibility of the spiral surface through phase tuning. In the long-range ordered regime, competing spiral and spin density wave orders emerge as a function of applied magnetic field, among which a possible order-by-disorder transition is identified.
pacs:
Introduction. A spiral spin-liquid (SSL) is an exotic type of correlated paramagnetic state where the low energy dynamics consist of collective spiral correlations Bergman et al. (2007); Lee and Balents (2008); Mulder et al. (2010); Zhang and Lamas (2013); Niggemann et al. (2020); Attig and Trebst (2017); Balla et al. (2020, 2019); Lee and Balents (2008); Chen (2017); Yao et al. (2021); Huang et al. (2022); Buessen et al. (2018); Liu et al. (2020); Mohylna et al. (2022). A characteristic feature of a SSL is that the propagation vectors of the degenerate spiral ground states form a continuous surface in reciprocal space Bergman et al. (2007). Such an unusual yet clearly defined feature has stimulated a strong interest by the community to experimentally identify and understand SSLs in real materials Tristan et al. (2005); T. Suzuki and H. Nagai and M. Nohara and H. Takagi (2007); Krimmel et al. (2009); Matsuda et al. (2010); MacDougall et al. (2011); Zaharko et al. (2011); Nair et al. (2014); MacDougall et al. (2016); Ge et al. (2017); Graham et al. (2023); Gao et al. (2017, 2020); Rosales et al. (2022); Guratinder et al. (2022); Chamorro et al. (2018); Bai et al. (2019); Tsurkan et al. (2021); Haraguchi et al. (2019); Abdeldaim et al. (2020); Otsuka et al. (2020); Wessler et al. (2020); Bordelon et al. (2021); Balz et al. (2016); Pohle et al. (2021); Takahashi et al. (2024); Gao et al. (2022a); Cole et al. (2023); Gao et al. (2022b); Hsieh et al. (2022). Through experimental and theoretical studies, compounds with a bipartite lattice, e.g. the honeycomb Balz et al. (2016); Pohle et al. (2021); Gao et al. (2022a); Cole et al. (2023) and diamond Graham et al. (2023); Balz et al. (2016); Pohle et al. (2021); Gao et al. (2017); Bai et al. (2019); Guratinder et al. (2022) lattices, have been demonstrated as the most fertile hosts of SSLs.
Figure 1: (a) AB-stacked triangular bilayers formed by the Fe3+ ions in Cs3Fe2Cl9. Atoms belonging to the neighboring bilayers are shown in red and blue, respectively. The $J_{1}$, $J_{2}$, and $J_{3}$ bonds are shown by thick black lines, thin black lines, and thin gray lines, respectively. Arrows indicate the spin directions of the collinear ground state with $\bm{q}=(\frac{1}{2},0,0)$. (b) The AB-stacked triangular bilayers viewed along the $c$ axis. (c) Refinement result of the powder neutron diffraction data measured on POWGEN at $T$ = 1.6 K. Data points are shown as red circles. The calculated pattern is shown as the black solid line. The vertical bars indicate the positions of the structural (upper) and magnetic (lower) Bragg peaks for Cs3Fe2Cl9. The blue line at the bottom shows the difference of measured and calculated intensities. The goodness-of-fit parameters are $R_{\textrm{p}}=18.6\%$ and $R_{\textrm{wp}}=10.2\%$.
Identifying novel SSL hosts is crucial for the realization of exotic spin textures like skyrmions Gao et al. (2020); Rosales et al. (2022); Takeda et al. (2024) and subdimensional quasiparticles like fractons Yan and Reuther (2022); Pretko et al. (2020), and will also establish new candidate compounds to study the thermal and quantum order-by-disorder (ObD) transitions that are elusive in real materials Villain et al. (1980); Henley (1989); Green et al. (2018); Bergman et al. (2007); Mulder et al. (2010). According to theoretical studies, SSLs can be classified by their codimension, a quantity that characterizes the dimensional difference between the spiral surface and the host system Yao et al. (2021). Experimentally identified SSLs, including those observed on the diamond and honeycomb lattices, exhibit either a two-dimensional (2D) spiral surface on a three-dimensional (3D) lattice Gao et al. (2017); Bai et al. (2019); Guratinder et al. (2022); Gao et al. (2022b); Graham et al. (2023) or a one-dimensional (1D) spiral surface, i.e. a degenerate line, on a 2D lattice Balz et al. (2016); Pohle et al. (2021); Takahashi et al. (2024); Gao et al. (2022a), thus all falling in the codimension one category. Although codimension-two SSLs have been predicted to exist on the AB-stacked triangular lattice Yao et al. (2021); Liu et al. (2016); Hoang and Diep (2012); Rastelli et al. (1983), their experimental realizations remain an open question.
Recent transport and magnetic characterizations of Cs3Fe2Cl9 unveil possible honeycomb physics on a 3D lattice Ishii et al. (2021). In this compound, magnetic Fe3+ ions with spin $S=\frac{5}{2}$ form AB-stacked triangular bilayers as shown in Figs. 1(a) and (b). Under dominant ferromagnetic (FM) interactions with exchange energy strength $J_{1}$ within the bilayers, a codimension-two SSL with a 1D spiral surface in the integer-$l$ planes may exist for $|J_{3}/J_{2}|>1/6$, where the threshold ratio is the same as that of the codimension-one SSL on the original honeycomb lattice Mulder et al. (2010). Although the magnetic ground state and spin dynamics in Cs3Fe2Cl9 remain unexplored, a rich phase diagram has been established Ishii et al. (2021), suggesting the existence of strong magnetic frustration in this compound.
Here, through elastic and inelastic neutron scattering experiments on both single crystal and polycrystalline samples, we show that a codimension-two SSL with a uniaxial anisotropy is realized in Cs3Fe2Cl9. The sprial surface’s visibility is phase tuned as a function of the wavevector normal to the honeycomb plane. By combining neutron diffraction experiments and classical Monte Carlo simulations, we clarify the eight field-induced ordered phases as competing spiral and spin density wave (SDW) orders, among which a possible order-by-disorder transition is identified.
Magnetic ground state. Powder neutron diffraction experiments were performed on POWGEN Huq et al. (2011) at the Spallation Neutron Source (SNS) of the Oak Ridge National Laboratory (ORNL) to determine the magnetic LRO in Cs3Fe2Cl9 below $T_{N}\sim 5.4$ K. Details on the sample preparation and neutron scattering experiments are presented in the Supplemental Material sup . As shown in Fig. 1(c), magnetic Bragg peaks belonging to the propagation vector $\bm{q}_{\textrm{I}}=(\frac{1}{2},0,0)$ are observed at low temperatures. Through Rietveld refinements Rodriguez-Carvajal (1993), the magnetic ground state is determined to be collinear as shown by arrows in Fig. 1(a), with an ordered moment size of 4.23(6) $\mu_{\textrm{B}}$. This magnetic order is similar to the magnetic ground state of the isostructural Cs3Fe2Br9 Brüning et al. (2021).
Spin Dynamics and Modeling. To determine the exchange coupling strengths, inelastic neutron scattering (INS) experiments were performed at CNCS Ehlers et al. (2011) at SNS of ORNL sup . Figure 2(a) presents the INS spectra collected at $T=2$ K with an incident neutron energy of $E_{i}=3.32$ meV. Even for the powder sample, it is clear that there are two highly dispersive magnon modes centered around $E\sim 0.6$ and 1.8 meV energy transfer. From Fig. 2(a) and also the $E_{i}=1.0$ meV data in the Supplemental Material sup , an excitation gap of $\Delta\sim 0.2$ meV is observed at wavevector transfer $Q\sim 0.6$-1, suggesting the existence of a uniaxial single-ion anisotropy (SIA) that stabilizes the collinear ground state.
Using linear spin wave theory as implemented in the SpinW program Toth and Lake (2015), $\chi^{2}$-fits to the INS spectra were performed to analyze the spin interactions. A minimal $J_{123}$-$D_{z}$ model with Hamiltonian $\mathcal{H}=\sum_{\langle ij\rangle\in n}J_{n}\bm{S}_{i}\cdot\bm{S}_{j}+D_{z}(S_{z})^{2}$ is employed in our calculations, which considers Heisenberg exchange interactions up to the third neighbors as shown in Fig. 1 plus a uniaxial SIA term, $D_{z}$. As compared in Fig. 2, the $J_{123}$-$D_{z}$ model with fitted coupling strengths of $J_{1}=-0.238(2)$, $J_{2}=0.085(2)$, $J_{3}=0.053(1)$, and $D_{z}=-0.037(2)$ meV reproduces the INS spectra. The dominant FM $J_{1}$ together with comparable strengths of $J_{2}$ and $J_{3}$ favor SSLs Mulder et al. (2010); Yao et al. (2021); Liu et al. (2016), while the relatively high magnitude of $|D_{z}|\sim 0.65J_{3}$ indicates the importance of the SIA.
Figure 2: (a) Inelastic neutron scattering spectra $S(q,\omega)$ for Cs3Fe2Cl9 powders at $T=2$ K measured on CNCS. (b) Simulated spectra for the $J_{123}$-$D_{z}$ model using linear spin wave theory.
Codimension-two SSLs. Figure 3(a) presents the diffuse neutron scattering pattern for Cs3Fe2Cl9 in the ($h$, $k$, 0) plane. Data were collected from CORELLI Ye et al. (2018) at SNS of ORNL using an $\sim 8$ mg crystal sup . At $T=6$ K, triangular shaped lobes are observed around the $K$-($\frac{1}{3}$, $\frac{1}{3}$, 0) points. As indicated by the dashed curves, the shape of the spiral surface is reproduced by a $J_{1}^{h}$-$J_{2}^{h}$ honeycomb-lattice model with $J_{2}^{h}/J_{1}^{h}=0.62$, where $J_{1}^{h}$ $(J_{2}^{h})$ is equivalent to $J_{2}$ ($J_{3}$) in the $J_{123}$-$D_{z}$ model. The equal-time spin correlations for the fitted $J_{123}$-$D_{z}$ model can be calculated using the self-consistent Gaussian approximation (SCGA) method Canals and Garanin (2001); sup . Since the critical correlations are underestimated in the SCGA method, a reduced $T=5$ K is assumed in the calculations to better describe the experimental data sup . As compared in Fig. 3(a), the calculated and experimental intensity distributions agree well with each other, both following the spiral surface of a SSL on a honeycomb lattice.
[…]
Competing orders and a possible ObD transition. In the absence of SIA, SSLs on a honeycomb lattice have been predicted to exhibit an ObD transition at low temperatures Bergman et al. (2007); Mulder et al. (2010); Liu et al. (2020). In the regime of $J_{2}^{h}/J_{1}^{h}>0.5$, the magnetic propagation vector $\bm{q}$ of the ObD phase lies at the corners of the triangular-shaped spiral surface Mulder et al. (2010); Liu et al. (2020). This is obviously not the case for Cs3Fe2Cl9 with $\bm{q}_{\textrm{I}}=(\frac{1}{2},0,0)$ as revealed in Fig. 1(c), suggesting the strong impact of the SIA.
However, the rich phase diagram reported for Cs3Fe2Cl9 Ishii et al. (2021), which is also reproduced in Fig. 4(a), indicates the the possibility of an ObD in magnetic field. In fields of $3\lesssim H\lesssim 6$ T, multiple phases, II-VI, emerge. Two additional phases exist at $H\gtrsim 6$ T, including a $\frac{1}{2}$-magnetization plateau phase, VII, up to $\sim 12$ T and a transitional phase, VIII, that precedes the field-polarized FM phase at $H\gtrsim 15$ T.
[…]
Due to the limited number of magnetic reflections collected on WAND2, we resort to classical Monte Carlo simulations to understand the rich phases in Cs3Fe2Cl9. As explained in the Supplemental Material sup , the fitted $J_{123}$-$D_{z}$ model exhibits a relatively simple phase diagram that contains only two main phases with $\bm{q}_{\textrm{I}}=(\frac{1}{2},0,0)$. Therefore, starting from the fitted $J_{123}$-$D_{z}$ model, we examined the impacts of weak perturbations from isotropic exchange interactions up to the tenth neighbors one-by-one. The exchange paths of each perturbations and the representative theoretical phase diagrams are presented in the Supplemental Material sup .
Following this method, we found that by adding a weak fifth-neighbor coupling $J_{5}=-0.008$ meV that only marginally impacts the INS spectra and diffuse scattering patterns sup , the calculated phase diagram, shown in Fig. 4(f), reproduces the main phases observed in experiments. Phases VI and VI′′ (Phase V) adjacent to phases I and VII exhibit similar diffraction pattern as that in phase VI (V). These primed phases, V, VI, and VI′′, are absent in experiments and may arise from finite size effects in our simulations. The magnetic orders in the main phases are described in Figs. 4(g)-(j), with their corresponding Ansatz and structure factor plots presented in the Supplemental Material sup . As is consistent with the experimental data, similar propagation vectors are observed in phases III-V. Among them, phases III and IV are determined to be of the spiral-type orders by inspecting the spin configuration snapshots in the Monte Carlo simulations, where the transition between phases III and IV can be attributed to the modulation of the $\bm{q}=0$ magnitude sup . As a contrast, both phases V and VI are revealed to be of the SDW orders with spins along the $c$ axis despite their different $\bm{q}$ vectors. In phases III-V where sufficient numbers of magnetic Bragg peaks can be extracted in our neutron diffraction experiments, the proposed magnetic structures agree with the experimental dataset sup , thus verifying the results of our Monte Carlo simulations.
Since $\bm{q}_{\textrm{VI}}$ in phase VI is the predicted propagation vector of the ObD theory Mulder et al. (2010); Liu et al. (2020), it is tantalizing to ascribe the transition from the SSL state to phase VI as entropy-driven. Compared to the theoretical phase diagram for the fitted $J_{123}$-$D_{z}$ model, it can be concluded that in the presence of a magnetic field, the $J_{5}$ perturbations favor orders with propagation vectors over the spiral surface, among which $\bm{q}_{\textrm{VI}}$ is stabilized by thermal fluctuations at higher temperatures. The proposed SDW character of phase VI, however, poses a challenge in accurately determining the free energy for states above $T_{N}$ under the current theoretical framework Bergman et al. (2007), making it difficult to exclude perturbations other than thermal fluctuations.
Conclusion. Our neutron scattering experiments establish Cs3Fe2Cl9 as a host of codimension-two SSLs, where the high codimension is manifested through the phase tuning of the visibility of the spiral surface. By combining neutron diffraction and classical Monte Carlo simulations, we clarify the rich field-induced phases as competing spiral and SDW orders, revealing Cs3Fe2Cl9 as a candidate compound to study the ObD transition in frustrated magnets.
Acknowledgements.
We acknowledge helpful discussions with Gang Chen, Xu-Ping Yao, and James Jun He. This work was supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division. This research used resources at the Spallation Neutron Source (SNS) and the High Flux Isotope Reactor (HFIR), both are DOE Office of Science User Facilities operated by the Oak Ridge National Laboratory (ORNL). The proposal numbers are IPTS-28988 (POWGEN), 29069 (CNCS), 31251 (CORELLI), and 29937 (WAND2). Numerical calculations in this paper were partly completed on the supercomputing system in the Supercomputing Center of USTC. Works at USTC were funded by the National Science Foundation of China (NSFC) under the Grant No. 12374152.
Supplemental Materials for:
Codimension-Two Spiral Spin-Liquid in the Effective Honeycomb-Lattice Compound Cs3Fe2Cl9
I Powder sample synthesis
Cs3Fe2Cl9 is extremely air sensitive and thus all handling of reagents and products was performed inside an inert atmosphere glovebox. Anhydrous FeCl3 (as-received) and CsCl (dried in air at 400 °C) were loaded into a fused silica crucible. The crucible was placed inside a fused silica ampoule, transferred to a vacuum line without exposure to air, purged with argon and sealed with approximately $\frac{1}{4}$ atm argon. The reagents were heated to 700 °C at a rate of 60 °C/h and held for 8 h prior to quenching in an ice-water bath. A boule was easily extracted from the silica ampoule, which was then ground and the powder was sealed with argon gas in a manner as done for the initial reaction. The powder was annealed at 250 °C for 10 d, with an external thermocouple utilized to verify the annealing temperature; the furnace was turned off and the sample was allowed to cool naturally.
II Single crystal growth
Single crystals of Cs3Fe2Cl9 were synthesized through a solvothermal route with a variety of concentrations and temperature profiles tested. One consistent feature of the growths was that a small excess of CsCl was found to improve the size and apparent quality of the crystals obtained. In the best growth conditions 1 mmol of anhydrous FeCl3 and 2 mmol anhydrous CsCl were weighed out in a glovebox under argon and removed in a small, closed vial. This vial was then emptied into a 23 mL Parr A280AC PTFE liner in air with concentrated HCl added immediately after, with additional HCl being used as a wash of the vial to ensure all material was transferred, with a total of 5 mL concentrated HCl being added to PTFE liner, which was subsequently sealed in a Parr 4749 general purpose acid digestion vessel. The vessel was placed in an oven which was ramped to 220 °C over one hour and held at this temperature for 6 hours to allow the vessel to come to temperature and allow the reactants to be fully dissolved in the concentrated HCl. The vessel was then cooled to 30 °C over 48 hours before the oven was turned off, with slower cooling rates not providing significant improvements in crystal size.
Due to the pressures involved and the porosity of PTFE the exterior of the liner was coated in a water-soluble green chloride salt due to the corrosion of the acid digestion vessel which was removed before the PTFE liners were opened. The liners were opened in air, with the resulting crystals metastable in the mother liquor, though undergoing a process of dissolution and recrystallization if stored under the mother liquor long-term. To avoid this, and because the crystals were highly air-sensitive they were immediately transferred into concentrated HCl under which they could be safely stored and additionally helped to remove the mother liquor from the crystals. The crystals stored under concentrated HCl were transferred to a glove bag which was filled with argon to create an air free environment safe from corrosion where the crystals were removed from the concentrated HCl, dried via filtration and transferred into a vial under argon which, free of HCl, was then transferred into a helium filled glovebox for long term storage. Crystals obtained through this method were up to 25 mg, though the largest unambiguously single crystal was 7.5 mg which was selected for neutron diffraction experiments on CORELLI and WAND2.
III Magnetization measurements
Magnetization measurements were performed on a single crystal weighing 4.5(1) mg using a Quantum Design Magnetic Property Measurement System MPMS-3. Due to the air sensitivity of the crystal it was coated in and secured to a quartz post with silicone vacuum grease. Magnetization measurements revealed similar phase transitions as previously reported Ishii et al. (2021) and were used to plan the neutron diffraction experiment.
Figure S1: Refinement result of the powder neutron diffraction data measured on POWGEN at $T$ = 15 K. Data points are shown as red circles. The calculated pattern is shown as the black solid line. The vertical bars indicate the positions of the structural Bragg peaks for Cs3Fe2Cl9. The blue line at the bottom shows the difference of measured and calculated intensities.
IV Neutron diffraction experiments on POWGEN
Powder neutron diffraction experiments were performed on POWGEN Huq et al. (2019) at the Spallation Neutron Source (SNS) of the Oak Ridge National Laboratory (ORNL). About 5 g powder of Cs3Fe2Cl9 was filled into an air-tight vanadium can in a helium filled glovebox. An orange cryostat was utilized to reach a base temperature of 2 K. Data reduction was performed using the MANTID software Arnold et al. (2014).
Figure S1 summarizes the refinement result of the neutron diffraction pattern collected at $T=15$ K. No secondary reflections are observed in the diffraction pattern, which confirms the phase purity of our sample. Refined crystal structure parameters are listed in Table S1. Due to the lacking of neutron diffraction data in the high-$Q$ region, we cannot reliable fit the thermal parameters. Therefore, a uniform thermal factor of $B_{\textrm{iso}}=0.1$2 is assumed for all atoms. Figure S2 plots the temperature evolution of the diffraction pattern. Below $T_{\textrm{N}}\sim 5.5$ K, magnetic Bragg peaks belonging to $\bm{q}=(\frac{1}{2},0,0)$ are observed. This transition temperature is consistent with that observed in magnetic susceptibility.
Figure S2: Temperature evolution of the powder neutron diffraction data for Cs3Fe2Cl9 collected on POWGEN. Positions of the representative magnetic Bragg peaks $(\frac{1}{2}00)$, $(\frac{1}{2}01)$, and $(\frac{1}{2}02)$ are indicated.
The refinement result of the neutron diffraction pattern collected at $T=1.6$ K is shown in Fig. 1 of the main text. Table S2 summarizes the refined crystal structure parameters. The refined magnitude of the ordered moment is 4.23(6) $\mu_{\textrm{B}}$.
Table S1: Refined crystal structure parameters for Cs3Fe2Cl9 at 15 K. The space group is $P6_{3}/mmc$ with lattice constants $a=7.1784(1)$ and $c=17.6374(3)$. A uniform thermal parameter of $B_{\textrm{iso}}=0.1$2 is assumed for all atoms. The goodness-of-fit parameters are $R_{p}=18.6\%$ and $R_{wp}=10.2\%$.
Atom $x$ $y$ $z$ site
Cs1 0 0 1/4 $2b$
Cs2 1/3 2/3 0.0826(2) $4f$
Fe 1/3 2/3 0.8458(1) $4f$
Cl1 0.519(4) 0.0330(3) 1/4 $12j$
Cl2 0.8248(2) 0.6496(4) 0.0892(1) $12k$
Table S2: Refined crystal structure parameters for Cs3Fe2Cl9 at 1.6 K. The space group is $P6_{3}/mmc$ with lattice constants $a=7.1784(1)$ and $c=17.6369(2)$. A uniform thermal parameter of $B_{\textrm{iso}}=0.1$2 is assumed for all atoms. The goodness-of-fit parameters are $R_{p}=15.2\%$ and $R_{wp}=11.3\%$.
Atom $x$ $y$ $z$ site
Cs1 0 0 1/4 $2b$
Cs2 1/3 2/3 0.0829(2) $4f$
Fe 1/3 2/3 0.8459(1) $4f$
Cl1 0.510(4) 0.0325(3) 1/4 $12j$
Cl2 0.8247(2) 0.6495(4) 0.0891(1) $12k$
V Inelastic neutron scattering experiments on CNCS
Inelastic neutron scattering (INS) experiments on powder sample of Cs3Fe2Cl9 were performed on CNCS Ehlers et al. (2011) at the SNS of the ORNL. About 5 g powder was sealed in an aluminum can in a helium filled glovebox. An orange cryostat was utilized to reach a base temperature of 5 K. Measurements were taken with incident neutron energies of $E_{i}=6.59$, 3.32, and 1.0 meV in the high flux chopper configuration at $T=2$ and 10 K. For each measuring condition, data were collected on an empty can and subtracted as background. Data reduction and projection were performed using the Mslice program in DAVE Azuah et al. (2009).
Figure S3: INS spectra for Cs3Fe2Cl9 measured on CNCS with an incident neutron energy of $E_{i}=1.0$ meV at $T=2$ K. Data were collected for approximately 8 hrs with the source operating at a power of 1.4 MW.
Figure S3 presents the INS spectra for Cs3Fe2Cl9 measured with an incident neutron nergy of $E_{i}=1.0$ meV at $T=2$ K. An excitation gap of $\sim 0.2$ meV is resolved, which is consistent with the existence of uniaxial SIA in Cs3Fe2Cl9.
VI Diffuse neutron scattering experiments on CORELLI
Single crystal diffuse neutron scattering experiments were performed on CORELLI Ye et al. (2018) at the SNS of the ORNL. A crystal (mass $\sim 7.5$ mg) was aligned with the $c$ axis vertical. A closed cycle refrigerator (CCR) was employed to reach temperatures $T$ down to 5 K. Data were acquired by rotating the sample in 1.5 steps, covering a total range of 360. The counting time at each rotation angle was approximately 1.5 mins with the source operating at a power of 1.4 MW. Data reduction and projection were performed using the MANTID software Arnold et al. (2014). Data collected at $T=20$ K were subtracted as background. Theoretical diffuse neutron scattering patterns were calculated using the self-consistent Gaussian approximation (SCGA) method Conlon and Chalker (2010) as implemented in JuliaSCGA Jul ; Gao et al. (2022).
Figure S4 compares the diffuse scattering patterns calculated using the SCGA method for the fitted $J_{123}$-$D_{z}$ model at $T=5$ and 6 K. As explained in the main text, the SCGA method is a mean-field-based theory and may underestimate the critical scattering in systems with relatively strong SIA. If we assume $T=6$ K as that in experiments, the calculated scattering pattern will exhibit less contrast at momentum transfers over and out of the spiral surface. Therefore, a reduced temperature of $T=5$ K is assumed in Fig. 3 of the main text to compensate for the underestimated critical scattering. On the other hand, the comparison between the calculations at $T=5$ and 6 K also reveals that the scattering intensity between the integer-$l$ planes is reduced at lower temperatures. This temperature dependence confirms thermal excitations as the origin of the interplanar scattering.
Figure S4: Comparison between the diffuse scattering patterns calculated using the SCGA method for the fitted $J_{123}$-$D_{z}$ model at $T=5$ K (a-f) and 6 K (g-l). Panels (a-f) is a reproduction of Fig. 3 in the main text.
Figure S5: (a) Calculated diffuse scattering pattern in the ($h,~{}k,~{}0$) plane using the fitted $J_{123}$-$D_{z}$ model at $T=6$ K. (b) Temperature evolution of the calculated spin correlation function $\langle\bm{S}\cdot\bm{S}\rangle$ along ($h$, $h$, 0) around (0.5, 0.5, 0). The integration area is outlined in panel (a) by a dashed rectangle. Black dots indicate the peak positions at the corresponding temperatures through Gaussian fits. (c) Temperature evolution of the calculated spin correlations $\langle S_{x}\cdot S_{x}\rangle$ (red triangles), $\langle S_{z}\cdot S_{z}\rangle$ (yellow circles), and their ratio $\langle S_{x}\cdot S_{x}\rangle/\langle S_{z}\cdot S_{z}\rangle$ (black dots). The integration area is outlined in panel (a) by a dashed-line hexagon. Red line over the $\langle S_{x}\cdot S_{x}\rangle/\langle S_{z}\cdot S_{z}\rangle$ data is the fitting curve to the Arrhenius law $A_{1}\exp(-\frac{T-T_{N}}{\Delta_{\textrm{SIA}}})+A_{2}$ with fitted parameters of $A_{1}=-0.52(1)$, $\Delta_{\textrm{SIA}}=4.03$ K, and $A_{2}=0.90(1)$.
VII Effects of the uniaxial anisotropy on the SSL
Classical Monte Carlo simulations for the fitted $J_{123}$-$D_{z}$ model were performed to study the effects of the SIA on the SSL in Cs3Fe2Cl9. As a reference, the diffuse scattering pattern at $T=6$ K calculated by the SCGA method is shown in Fig. S5. Through classical Monte Carlo simulations that incorporate critical scattering near $T_{N}$, Fig. S5(b) summarizes the temperature evolution of the calculated scattering intensity along ($h+\frac{1}{2}$, $h+\frac{1}{2}$, 0), where the integration range is outlined by dashed rectangle in Fig. S5(b). For the fitted $J_{123}$-$D_{z}$ model, $T_{N}$ is found to be $\sim 5.5$ K. The separated peaks at $T/T_{N}>1$ correspond to thermally stabilized spiral correlations in the SSL state, while their gradually reduced separation with decreasing $T$ reveals the second-order character of the transition into the LRO. Figure S5(d) compares the in-plane ($x$) and out-of-plane ($z$) components of the calculated spin correlations as a function of $T$. Due to the existence of uniaxial SIA, the evolution of $\langle S_{x}\cdot S_{x}\rangle$ is nonmonotonic, and the ratio between the in-plane and out-of-plane components exhibits an Arrhenius-like behavior in a wide temperature regime.
VIII Neutron diffraction experiments on WAND2
Single crystal neutron diffraction measurements were performed using the WAND2 diffractometer Frontzek et al. (2018) at the High Flux Isotope Reactor HFIR with $\lambda=1.486$. A crystal with mass of $\sim 7.5$ mg was aligned with the $c$ axis vertical, and then sealed in an aluminum can with helium exchange gas. A vertical field cryomagnet was used, providing a base temperature of 1.5 K and a maximum field of $H=6$ T. Measurements were performed in phase I at $T$ = 1.5 K and no applied field rotating through 180° in 0.1° steps over 20 h. Measurements in phase II were made at $T$ = 1.5 K and $H=3.2$ T by rotating through 180° in 0.1° steps over 20 h. In phase III at $T=1.5$ K at $H=4.5$ T the sample was rotated through 90° in 0.1° steps over 10 h. In phase IV, at $T=1.5$ K and $H=5.5$ T the sample was rotated through 90° in 0.1° steps over 10 h. In phase V, at $T=4$ K and $H=5.5$ T the sample was rotated through 90° in 0.1° steps over 10 h. In phase VI, at $T=5.05$ K and $H=5.5$ T the sample was rotated over 90° in 0.1° steps with a total measurement time of 26 h.
Figure S6: Single crystal neutron diffraction pattern for Cs3Fe2Cl9 in phases (a) III, (b) IV, and (c) V in the ($h$,$k$,0) plane. The experimental conditions are listed in the text. The dashed rectangle emphasizes the area in which the relative intensity of the $\bm{q}_{\textrm{III}}$ and $2\bm{q}_{\textrm{III}}$ reflections are compared in the text.
Figure S6 compares the neutron diffraction patterns in phases III, IV, and V. As explained in the main text, similar diffraction patterns were observed in these three phases. Especially, their magnetic propagation vector stays at the same $\bm{q}_{\textrm{III}}=\frac{5}{6}\times(\frac{1}{3},\frac{1}{3},0)$, with weak $2\bm{q}_{\textrm{III}}$ higher harmonics observed along the Brillouin zone boundary. Despite their similarities, careful comparisons of the diffraction patterns reveal some differences among these three phases. As emphasized by the dashed rectangle in each panel, the intensity of the $2\bm{q}_{\textrm{III}}$ reflection in phase III is weaker than that in phase IV, which suggests higher field-induced magnetization in phase IV. Comparing phases IV and V, the $\bm{q}_{\textrm{III}}$ reflection becomes weaker in phase V while the $2\bm{q}_{\textrm{III}}$ reflection stays almost unchanged. This evolution is consistent with a transition from a spiral structure in phase IV to a spin density wave (SDW) structure in phase V.
IX Refinements of the nuclear and magnetic structures
Figure S7 compares the experimental intensity of the nuclear Bragg reflections observed on WAND2 to the calculations assuming the same crystal structure as that listed in Table S2. The goodness-of-fit parameters are $R_{F2}=26.9\%$ and $R_{F2w}=22.4\%$. The agreement between the observed and calculated intensity confirms the good quality of our crystal sample.
Figure S7: Comparison between the calculated and observed intensity of the nuclear reflections collected on WAND2.
Figure S8: Comparison between the calculated and observed intensity of the magnetic reflections collected on WAND2 for phase I (a), III (b), IV (c), and V (d).
Figure S8 summarizes the fitting results of the magnetic reflections collected on WAND2 in phases I (a), III (b), VI (c), and V (d). The measuring conditions are listed in the main text. In phase VI, only two nonequivalent magnetic reflections were obtained for each magnetic domain, which does not allow a reliable analysis of the magnetic structure. For phase I, the collinear magnetic structure as shown in Fig. 1(a) of the main text was assumed in the fits. The fitted magnitude of the ordered magnetic moment is 2.9(2) $\mu\rm{{}_{B}}$, with goodness-of-fit parameters $R_{F2}=20.3\%$ and $R_{F2w}=24.4\%$. For phases III-V, a helical magnetic structure was assumed based on our classical Monte Carlo simulations. The magnitudes of the in-plane magnetic moment, $M_{\parallel}$, and out-of-plane magnetic moment, $M_{\perp}$, were treated as fitting parameters. For phase III, the fitted magnitudes are $M_{\parallel}=1.9\pm 1.5$ and $M_{\perp}=1.9\pm 0.4$ $\mu\rm{{}_{B}}$, with goodness-of-fit parameters $R_{F2}=13.0\%$ and $R_{F2w}=14.7\%$. For phase IV, the fitted magnitudes are $M_{\parallel}=2.1\pm 1.6$ and $M_{\perp}=1.8\pm 0.5$ $\mu\rm{{}_{B}}$, with $R_{F2}=9.0\%$ and $R_{F2w}=10.1\%$. For phase V, the fitted magnitudes are $M_{\parallel}=1.8\pm 4.1$ and $M_{\perp}=1.2\pm 0.8$ $\mu\rm{{}_{B}}$, with $R_{F2}=3.2\%$ and $R_{F2w}=3.7\%$. Considering the standard deviations, the fitted magnitude of the in-plane moment $M_{\parallel}$ is consistent with our classical Monte Carlo simulations, where phases III and IV exhibit an elongated helical structure, and phase V exhibits a spin density wave structure with only the $S_{z}$ components.
X Classical Monte Carlo simulations
Classical Monte Carlo simulations for the $J_{123}$-$D_{z}$ model
$\mathcal{H}=\sum_{\langle ij\rangle\in n}J_{n}\bm{S}_{i}\cdot\bm{S}_{j}+D_{z}(S_{z})^{2}$ (S1)
and its perturbed derivatives were performed using the SpinMC code that implements the single spin flip Metropolis algorithm bue . Unless alternately specified, a $12\times 12\times 4$ supercell with 2304 spins was employed in simulations. Observables including the heat capacity and magnetization were averaged over $10^{6}$ of measurement sweeps after $3\times 10^{5}$ thermalization sweeps, where each sweep represents 2304 attempted spin flips at randomly selected sites. Magnetic structure factors were calculated through fast Fourier transform using the FFTW package fft . The parallel tempering algorithm was utilized to facilitate thermal equilibrium, which was performed simultaneously over 128 replicas on 128 cores with a geometric series of temperatures between 1 and 10 K. After every 10 Monte Carlo sweeps, a replica exchange was attempted and the successful exchange rates are above $\sim 25\%$ for all the neighboring replicas.
X.1 The $J_{123}$ model
Figure S9 summarizes the simulation results for the $J_{123}$ model with no SIA. The coupling strengths are $J_{1}=-0.238$, $J_{2}=0.085$, $J_{3}=0.053$ meV as fitted from the INS experimental data. In zero field, the $J_{123}$ model exhibits a long-range order transition at $T_{N}\sim 3.5$ K, which is lower than that observed in experiments. The phase diagram shown in Fig. S9(a) mainly consists of two phases. Magnetic structure factors shown in Figs. S9(b-d) with $\bm{q}\sim\frac{4}{3}\times(\frac{1}{3},\frac{1}{3},0)$ in low magnetic fields and $\bm{q}=(\frac{1}{2},\frac{1}{2},0)$ in high magnetic fields, which is very different from the experimental observations.
Figure S9: (a) Theoretical phase diagram for the $J_{123}$ model calculated by the classical Monte Carlo simulations. Pseudocolor corresponds to the value of heat capacity in arbitrary units (arb. units). (b-d) Magnetic structure factors $\langle\bm{S}\cdot\bm{S}\rangle$ calculated at $T=1$ K in a field of (b) 0 T, (c) 5 T, and (d) 10 T. Calculations of the magnetic structure factor were performed on a $18\times 18\times 4$ supercell with $3\times 10^{5}$ thermalization sweeps and $5\times 10^{5}$ measurement sweeps.
Figure S10: (a) Theoretical phase diagram for the $J_{123}+D_{z}$ model calculated by the classical Monte Carlo simulations. An easy-axis SIA of $D_{z}=-0.037$ meV is added to the $J_{123}$ model considered in Fig. S9. Pseudocolor corresponds to the value of heat capacity in arbitrary units (arb. units). (b-d) Magnetic structure factors $\langle\bm{S}\cdot\bm{S}\rangle$ calculated at $T=1$ K in a field of (b) 0 T, (c) 6 T, and (d) 10 T.
Figure S11: Exchange paths for the perturbational couplings $J_{4}$ $J_{10}$ that are considered in our Monte Carlo simulations.
X.2 The $J_{123}+D_{z}$ model
Figure S10 summarizes the simulation results for the $J_{123}+{D_{z}}$ model that considers isotropic exchange interactions up to the third neighbors and a uniaxial SIA. The coupling strengths are $J_{1}=-0.238$, $J_{2}=0.085$, $J_{3}=0.053$, and $D_{z}=-0.037$ meV as fitted from the INS experimental data. In zero field, a magnetic long-range order transition is observed at $T_{N}\sim 5.8$ K, which is close to the experimental observation. Although the phase diagram shown in Fig. S10(a) mainly consists of two phases as that in Fig. S9(a), magnetic structure factors shown in Figs. S10(b-d) reveal that the propagation vector $\bm{q}$ stays at $(\frac{1}{2},\frac{1}{2},0)$ r.l.u. in both phases. This calculation indicates that the experimentally observed $\bm{q}_{\textrm{I}}=(\frac{1}{2},\frac{1}{2},0)$ r.l.u. in phase I is due to the existence of uniaxial SIA, yet the stabilization of the field-induced phases III, IV, V, and VI requires further perturbations in the spin Hamiltonian.
Figure S12: (a) Theoretical phase diagram for the $J_{123}+D_{z}+J_{6}$ model calculated by the classical Monte Carlo simulations. A perturbative sixth-neighbor interaction of $J_{6}=0.01$ meV is added to the $J_{123}+D_{z}$ model considered in Fig. S10. Pseudocolor corresponds to the value of heat capacity in arbitrary units (arb. units). (b-d) Magnetic structure factors $\langle\bm{S}\cdot\bm{S}\rangle$ calculated at $T=1$ K in a field of (b) 0 T, (c) 5 T, and (d) 10 T.
X.3 An overview of the perturbation effects
As the theoretical phase diagram for the $J_{123}+D_{z}$ model cannot reproduce the experimental results, extensive classical Monte Carlo simulations were performed to explore possible perturbative terms that may stabilize the experimentally observed additional phases in intermediate fields. By confining the perturbations to a single-term Heisenberg exchange interaction up to the tenth neighbors, we calculated the phase diagram and magnetic structure factors for each perturbed model. The exchange paths for $J_{4}$-$J_{10}$ are indicated in Fig. S11.
Figure S13: (a) Theoretical phase diagram for the $J_{123}+D_{z}+J_{5}$ model calculated by the classical Monte Carlo simulations. A perturbative fifth-neighbor interaction of $J_{5}=-0.008$ meV is added to the $J_{123}-D_{z}$ model considered in Fig. S10. Pseudocolor corresponds to the value of heat capacity in arbitrary units (arb. units). (b-d) Magnetic structure factors $\langle\bm{S}\cdot\bm{S}\rangle$ calculated at $T=1$ K in a field of (b) 0 T, (c) 5 T, and (d) 10 T. Calculations of the magnetic structure factor were performed on a $18\times 18\times 4$ supercell with $3\times 10^{5}$ thermalization sweeps and $5\times 10^{5}$ measurement sweeps. White arrows indicate the positions of weak magnetic Bragg peaks.
Figure S14: Theoretical $M/H$ curves calculated by the classical Monte Carlo simulations for the $J_{123}+D_{z}+J_{5}$ model. The strength of the applied field for each curve is indicated on the left. Data are successively shifted by 0.003 arb. units along the $y$ axis for clarity. The dashed line at $T=1.8$ K marks the lower limit of the measuring temperature in Ref. Ishii et al. (2021).
Through our calculations, it was found that all of the considered perturbations with a strength lower than $\sim 0.01$ meV are able to introduce additional phases in the intermediate field regime, yet their effects on the magnetic structure factors categorize them into two groups. The first group contains antiferromagnetic (AFM) $J_{6}$ and $J_{7}$, or ferromagnetic (FM) $J_{8}$, $J_{9}$, and $J_{10}$ perturbations. However, in zero field, these types of perturbations do not produce the experimentally observed $\bm{q}_{\textrm{I}}=(\frac{1}{2},\frac{1}{2},0)$ in phase I and thus contradicts the experimental observations. The second group contains AFM $J_{4}$ or FM $J_{5}$ perturbations. This group of perturbations stabilize $\bm{q}_{\textrm{I}}=(\frac{1}{2},\frac{1}{2},0)$ in zero field while introducing additional phases similar to that observed in experiments. For both groups of perturbations, a sign reversal eliminates the field-induced phases, leading to phase diagrams similar to that of the unperturbed model shown in Fig. S10(a).
Figure S15: Theoretical heat capacity $C_{p}$ calculated by the classical Monte Carlo simulations for the $J_{123}+D_{z}+J_{5}$ model in the field regime of (a) [0, 3], (b) [4, 7], and (c) [8, 14] T. Data are normalized by the maximal $C_{p}$ value, $C_{p}^{\textrm{max}}$. The strength of the applied field for each curve is indicated on the left. In each panel, data are successively shifted by 1 unit along the $y$ axis for clarity. The dashed line at $T=2.0$ K marks the lower limit of the measuring temperature in Ref. Ishii et al. (2021).
X.4 The $J_{123}+D_{z}+J_{6}$ model
As an illustration for the first group of perturbations, Fig. S10 summarizes the simulation results for the $J_{123}+{D_{z}}+J_{6}$ model that considers a sixth-neighbor AFM perturbation of $J_{6}=0.01$ meV. Comparison of the phase diagrams shown in Fig. S10(a) and Fig. S12(a) reveals that the perturbation of $J_{6}$ causes additional phases in the intermediate field regime between $\sim 4.5$ and 6.5 T. However, as shown in Fig. S12(b-d), the propagation vector of the long-range ordered phase in zero field deviates from $\bm{q}_{\textrm{I}}=(\frac{1}{2},\frac{1}{2},0)$ r.l.u. and thus fails to reproduce the experimental results. Similar effects are observed for the remaining perturbations in the first group, thus excluding them as the main perturbation to the $J_{123}+{D_{z}}$ Hamiltonian.
X.5 The $J_{123}+D_{z}+J_{5}$ model
As discussed in the previous section, AFM $J_{4}$ and FM $J_{5}$ belong to the second group of perturbations and have similar effects on the phase diagram. However, there is one slight difference for these two perturbations: In zero field, AFM $J_{4}$ and FM $J_{5}$ raises and lowers $T_{N}$, respectively. This observation favors FM $J_{5}$ since the experimental $T_{N}\sim 5.4$ K is lower than the theoretically predicted $T_{N}\sim 5.8$ K for the unperturbed $J_{123}+D_{z}$ model.
Figures S13-S15 summarize the simulation results for the $J_{123}+{D_{z}}+J_{5}$ model that considers a FM perturbation of $J_{5}=-0.008$ meV. This strength of $J_{5}$ is found to best describe the experimental phase diagram on a $12\times 12\times 4$ supercell. By carefully analyzing the field and temperature dependence of the magnetic susceptibility (see Fig. S14) and specific heat (see Fig. S15), we obtain the phase diagram shown in Fig. S13(a). The magnetic structure factors shown in Figs. S13(b-d) reproduce the $\bm{q}_{\textrm{I}}=(\frac{1}{2},\frac{1}{2},0)$ in zero field and $\bm{q}_{\textrm{III}}=\frac{5}{6}\times(\frac{1}{3},\frac{1}{3},0)$ at $H=5$ T. Therefore, we conclude that FM $J_{5}$ is the main perturbative term to the $J_{123}+{D_{z}}$ Hamiltonian, leading to a $J_{123}+{D_{z}}+J_{5}$ model that is employed in the main text.
Figure S16: Theoretical magnetic structure factors (a) $\langle S_{z}S_{z}\rangle$ and (b) $\langle S_{x}S_{x}\rangle$ in phase III for the $J_{123}+D_{z}+J_{5}$ model. Calculations were performed on a $15\times 15\times 4$ supercell using the classical Monte Carlo simulations. The strength of $J_{5}$ is slightly increased to $-0.01$ meV to compensate for the size effect so that the phase boundaries are similar to that shown in Fig. 4(f) of the main text. The temperature and magnetic field are set to 1 K and 4 T, respectively. Calculations were performed over $5\times 10^{5}$ measuring sweeps after $5\times 10^{5}$ thermalization sweeps. (c-h) Similar magnetic structure factors in phases IV, V, and VI. The calculation parameters are $T=1$ K and $H=5.5$ T in phase IV, $T=3.1$ K and $H=4.0$ T in phase V, and $T=5.0$ K and $H=4.0$ T in phase VI.
X.5.1 Spiral-type orders in phases III and IV
Comparisons of the direction-resolved magnetic structure factors and real-space spin configurations unveil the magnetic structures in each of the field-induced phases. Figure S16 compares the $\langle S_{z}S_{z}\rangle$ and $\langle S_{x}S_{x}\rangle$ components of the magnetic structure factors calculated for phases III, IV, V, and VI through the classical Monte Carlo simulations. In phases III and IV, both the $\langle S_{z}S_{z}\rangle$ and $\langle S_{x}S_{x}\rangle$ components exhibit long-range order, suggesting a spiral-type magnetic order that involves both in-plane and out-of-plane spin components. The magnetic structure factor along the $z$ direction is about 10 times higher than that along the $x$ direction, which indicates that the spiral order in phases III and IV is elongated along the $c$ axis due to the uniaxial SIA. Compared to that in phase III, the relatively stronger $2\bm{q}_{\textrm{III}}$ reflections in phase IV suggests stronger squaring-up effects in a higher magnetic field. This field dependence of the $2\bm{q}_{\textrm{III}}$ reflections is consistent with the experimental observations shown in Fig. S6.
X.5.2 SDW orders in phases V and VI
As shown in Fig. S16(f) and (h), the $\langle S_{x}S_{x}\rangle$ component of the magnetic structure factors in phases V and VI exhibit diffuse patterns, while sharp Bragg peaks are observed in the $\langle S_{z}S_{z}\rangle$ component. This observation indicates that magnetic moments in phases V and VI are ordered only along the $c$ axis, leading to sinusoidally modulated SDW orders in these two phases. We also note that comparison of the $\langle S_{z}S_{z}\rangle$ component in phases III (Fig. S16(a)), IV (Fig. S16(c)), and V (Fig. S16(e)) reveals a similar magnetic propagation vector $\bm{q}_{\textrm{III}}$ and its high harmonics $2\bm{q}_{\textrm{III}}$, which reproduces the experimental observations shown in Fig. S16.
XI Analytical expressions for the field-induced phases
To further verify the magnetic structures of the field-induced phases, we calculate the magnetic structure factors for the following ansatz on a honeycomb lattice to describe the spiral-type orders in phases III and IV (Eqn. (S2)) and the SDW orders in phases V and VI (Eqn. (S3)):
$\bm{M}(\bm{r})=\left\{\begin{array}[]{l}M_{\perp}\cos(\bm{q}\cdot\bm{r})\bm{n}_{1}+M_{\parallel}\sin(\bm{q}\cdot\bm{r})\bm{n}_{2}+M_{z}\bm{n}_{1}\textrm{,\qquad\qquad\ \ if}\ \bm{r}\in\{r_{1}\}\\ M_{\perp}\cos(\bm{q}\cdot\bm{r}+\phi)\bm{n}_{1}+M_{\parallel}\sin(\bm{q}\cdot\bm{r}+\phi)\bm{n}_{2}+M_{z}\bm{n}_{1}\textrm{,\quad if}\ \bm{r}\in\{r_{2}\}\end{array}\right.$ (S2)
$\bm{M}(\bm{r})=\left\{\begin{array}[]{l}M_{\perp}\cos(\bm{q}\cdot\bm{r})\bm{n}_{1}+M_{z}\bm{n}_{1}\textrm{,\qquad\ \ if}\ \bm{r}\in\{r_{1}\}\\ M_{\perp}\cos(\bm{q}\cdot\bm{r}+\phi)\bm{n}_{1}+M_{z}\bm{n}_{1}\textrm{,\quad if}\ \bm{r}\in\{r_{2}\}\end{array}\right.$ (S3)
Figure S17: (a) Theoretical magnetic structure factors (a) $\langle S_{z}S_{z}\rangle$ and (b) $\langle S_{x}S_{x}\rangle$ using the ansatz in Eqn. (S2) for phase III. Calculations were performed for a two-dimensional $30\times 30$ honeycomb superlattice through Fourier transform. (c-d) Magnetic structure factors using the ansatz in Eqn. (S2) for phases IV. (e) Magnetic structure factor $\langle S_{z}S_{z}\rangle$ using the ansatz in Eqn. (S3) for phase V. (d) Magnetic structure factor $\langle S_{z}S_{z}\rangle$ using the ansatz in Eqn. (S3) for phase VI.
In these expressions, $M_{\perp}$ and $M_{\parallel}$ are the ordered moments perpendicular to and parallel with the $ab$ plane, respectively. $M_{z}$ is the field induced moment along the $c$ axis. $\bm{n}_{1}$ is a unit vector along the $c$ axis, $\bm{n}_{2}$ is defined in a way that the vectors $\bm{q}$, $\bm{n}_{2}$, and $\bm{n}_{1}$ form a cartesian coordinate system. An additional phase factor $\phi$ is introduced for spins on the second sublattice ($\bm{r}\in\{r_{2}\}$).
Figure S17 summarizes the magnetic structure factors for each ansatz. $\bm{M}(\bm{r})$ in real space was first calculated on a $30\times 30$ superlattice, renormalized to equal moment size if $M_{z}\neq 0$, and then Fourier transferred to reciprocal space. The parameters are $\bm{q}=\bm{q}_{\textrm{III}}=\frac{5}{6}\times(\frac{1}{3},\frac{1}{3})$, $M_{\perp}=0.9$, $M_{\parallel}=0.1$, $\phi=\pi$, and $M_{z}=0.38$ (0.58) for phase III (IV). The parameters for phase V are $\bm{q}=\bm{q_{\textrm{III}}}=\frac{5}{6}\times(\frac{1}{3},\frac{1}{3})$, $M_{\perp}=1$, $M_{z}=0.48$, and $\phi=\pi$. The parameters for phase VI are $\bm{q}_{\textrm{VI}}=\frac{11}{9}\times(\frac{1}{3},\frac{1}{3},0)$, $M_{\perp}=0.5$, $M_{z}=0$, and $\phi=0$. By comparing the magnetic structure factors in Fig. S17 with those in Fig. S16, Fig. S6, and Fig. 4 in the main text, it is confirmed that the double peaks along over the Brillouin zone boundaries are due to high harmonics $2\bm{q}_{\textrm{III}}$ reflections induced by magnetic field, thus corroborating the proposed spiral-type and SDW orders in the field-induced phases.
XII Magnetic order in the magnetization plateau phase
A few possible magnetic orders have been proposed for the 1/2-magnetization plateau phase Ishii et al. (2021). Through classical Monte Carlo simulations, the order in the 1/2-magnetization plateau phase of the $J_{123}+{D_{z}}+J_{5}$ model is presented in Fig. S18. Further neutron diffraction experiments in fields above 6 T will be required to verify this magnetic order.
Figure S18: Magnetic order in phase VII viewed along the $c$ axis as determined from classical Monte Carlo simulations for the $J_{123}+{D_{z}}+J_{5}$ model. The bottom left (top right) part depicts the spin configuration for two sublattices (one sublattice).
XIII Impacts of the $J_{5}$ perturbation on the spin dynamics
Figure S19 presents the calculated INS spectra and diffuse neutron scattering pattern in the $(h,k,0)$ plane for the $J_{123}+D_{z}+J_{5}$ model with a perturbative $J_{5}=-0.008$ meV. The INS spectra are calculated by the linear spin wave theory. The diffuse neutron scattering pattern is calculated by the SCGA method at $T=5$ K. The calculated results only exhibit marginal difference compared to those of the $J_{123}+D_{z}$ model as presented in the main text.
Figure S19: (a) Calculated INS spectra and (b) diffuse neutron scattering pattern in the $(h,k,0)$ plane for the $J_{123}+D_{z}+J_{5}$ model with $J_{5}=-0.008$ meV. The diffuse neutron scattering pattern is calculated by the SCGA method at $T=5$ K.

### v2

@@
 A codimension-two spiral spin-liquid is a correlated paramagnetic state with one-dimensional ground state degeneracy hosted within a three-dimensional lattice.
-Here, via neutron scattering experiments and numerical simulations, we establish the existence of a codimension-two spiral spin-liquid in the effective honeycomb-lattice compound Cs3Fe2Cl9 and demonstrate the selective visibility of the spiral surface through phase tuning.
+Here, via neutron scattering experiments and numerical simulations, we establish the existence of a codimension-two spiral spin-liquid in the effective honeycomb-lattice compound Cs3Fe2Cl9, which demonstrates a novel path to spiral spin-liquids by overcoming the long-standing impediment of weak further-neighbor interactions.
 In the long-range ordered regime, competing spiral and spin density wave orders emerge as a function of applied magnetic field, among which a possible order-by-disorder transition is identified.
-pacs:
 Introduction.
@@
 (2022b); Hsieh et al.
+(2022); Wan et al.
+(2024); Andriushin et al.
 (2025).
@@
 (2019); Guratinder et al.
-(2022) lattices, have been demonstrated as the most fertile hosts of SSLs.
+(2022) lattices, have been demonstrated as the most fertile hosts of SSLs, where spiral correlations arise from the competition between the inter- and intra-sublattice interactions.
 Figure 1: (a) AB-stacked triangular bilayers formed by the Fe3+ ions in Cs3Fe2Cl9.
@@
 The $J_{1}$, $J_{2}$, and $J_{3}$ bonds are shown by thick black lines, thin black lines, and thin gray lines, respectively.
+The $J_{4}$ and $J_{5}$ bonds are indicated by red curved lines.
 Arrows indicate the spin directions of the collinear ground state with $\bm{q}=(\frac{1}{2},0,0)$.
@@
 The goodness-of-fit parameters are $R_{\textrm{p}}=18.6\%$ and $R_{\textrm{wp}}=10.2\%$.
+Figure 2: (a,c,e,g) Inelastic neutron scattering spectra $S(q,\omega)$ for Cs3Fe2Cl9 powders measured on CNCS with $E_{i}=3.32$ meV at $T=2$ (a) and 10 K (c), and with $E_{i}=1.0$ meV at $T=2$ (e) and 10 K (g).
+(b,f) Simulated spectra for the fitted $J_{1\textrm{-}5}$-$D_{z}$ model using linear spin wave theory (LSWT).
+The spectra in (b) and (f) are convolved with the instrumental energy resolution for $E_{i}=3.32$ and 1.0 meV, respectively.
+(d,h) Simulated spectra for the fitted $J_{1\textrm{-}5}$-$D_{z}$ model using the Landau-Langevin-Gilbert (LLG) method calculated at $T=10$ K.
 Identifying novel SSL hosts is crucial for the realization of exotic spin textures like skyrmions Gao et al.
@@
 (1983), their experimental realizations remain an open question.
+If proven true, this novel approach to spiral spin-liquids, where the intra-sublattice coupling arises from the relatively strong intralayer interactions, will relieve the long-standing obstacle of insufficient frustration on the original honeycomb lattice Gao et al.
+(2022a).
 Recent transport and magnetic characterizations of Cs3Fe2Cl9 unveil possible honeycomb physics on a 3D lattice Ishii et al.
@@
 (2011) at SNS of ORNL sup .
-Figure 2(a) presents the INS spectra collected at $T=2$ K with an incident neutron energy of $E_{i}=3.32$ meV.
+Figure 2(a) presents the INS spectra collected with an incident neutron energy of $E_{i}=3.32$ (a,c) and 1.0 meV (e,g) at temperatures $T=2$ (a,e) and 10 K (c,g), which are below and above $T_{N}$, respectively.
 Even for the powder sample, it is clear that there are two highly dispersive magnon modes centered around $E\sim 0.6$ and 1.8 meV energy transfer.
 From Fig.
-2(a) and also the $E_{i}=1.0$ meV data in the Supplemental Material sup , an excitation gap of $\Delta\sim 0.2$ meV is observed at wavevector transfer $Q\sim 0.6$-1, suggesting the existence of a uniaxial single-ion anisotropy (SIA) that stabilizes the collinear ground state.
+2(e), an excitation gap of $\Delta\sim 0.2$ meV is observed at wavevector transfer $Q\sim 0.6$ Å-1, suggesting the existence of a uniaxial single-ion anisotropy (SIA) that stabilizes the collinear ground state.
 Using linear spin wave theory as implemented in the SpinW program Toth and Lake (2015), $\chi^{2}$-fits to the INS spectra were performed to analyze the spin interactions.
-A minimal $J_{123}$-$D_{z}$ model with Hamiltonian $\mathcal{H}=\sum_{\langle ij\rangle\in n}J_{n}\bm{S}_{i}\cdot\bm{S}_{j}+D_{z}(S_{z})^{2}$ is employed in our calculations, which considers Heisenberg exchange interactions up to the third neighbors as shown in Fig.
+As explained in the Supplemental Material sup , a minimal $J_{1\textrm{-}5}$-$D_{z}$ model with Hamiltonian $\mathcal{H}=\sum_{\langle ij\rangle\in n}J_{n}\bm{S}_{i}\cdot\bm{S}_{j}+D_{z}(S_{z})^{2}$ is employed in our calculations, which considers Heisenberg exchange interactions up to the fifth neighbors as shown in Fig.
 1 plus a uniaxial SIA term, $D_{z}$.
 As compared in Fig.
-2, the $J_{123}$-$D_{z}$ model with fitted coupling strengths of $J_{1}=-0.238(2)$, $J_{2}=0.085(2)$, $J_{3}=0.053(1)$, and $D_{z}=-0.037(2)$ meV reproduces the INS spectra.
+2(a,b,e,f), the $J_{1\textrm{-}5}$-$D_{z}$ model with fitted coupling strengths of $J_{1}=-0.231(2)$, $J_{2}=0.082(2)$, $J_{3}=0.059(1)$, $J_{4}=0.0050(1)$, $J_{5}=0.0013(1)$ and $D_{z}=-0.032(2)$ meV reproduces the INS spectra.
 The dominant FM $J_{1}$ together with comparable strengths of $J_{2}$ and $J_{3}$ favor SSLs Mulder et al.
@@
 (2021); Liu et al.
-(2016), while the relatively high magnitude of $|D_{z}|\sim 0.65J_{3}$ indicates the importance of the SIA.
-Figure 2: (a) Inelastic neutron scattering spectra $S(q,\omega)$ for Cs3Fe2Cl9 powders at $T=2$ K measured on CNCS. (b) Simulated spectra for the $J_{123}$-$D_{z}$ model using linear spin wave theory.
+(2016), while the relatively high magnitude of $|D_{z}|\sim 0.54J_{3}$ indicates the importance of the SIA in stabilizing the ground state.
+The accuracy of the fitted model is further verified through the comparison of the INS spectra at $T=10$ K shown in Fig.
+2(c,d,g,h), where the calculations were performed using the Landau-Lifshitz method as implemented in the JuliaMD program Jul .
+At elevated temperatures, the magnon mode at higher energy becomes broadened and softened, and the magnon band at lower energy collapses towards the elastic line, which indicates the emergence of strong spin correlation in the paramagnetic regime that can be directly detected through diffuse neutron scattering.
 Codimension-two SSLs.
@@
 (2018) at SNS of ORNL using an $\sim 8$ mg crystal sup .
-At $T=6$ K, triangular shaped lobes are observed around the $K$-($\frac{1}{3}$, $\frac{1}{3}$, 0) points.
-As indicated by the dashed curves, the shape of the spiral surface is reproduced by a $J_{1}^{h}$-$J_{2}^{h}$ honeycomb-lattice model with $J_{2}^{h}/J_{1}^{h}=0.62$, where $J_{1}^{h}$ $(J_{2}^{h})$ is equivalent to $J_{2}$ ($J_{3}$) in the $J_{123}$-$D_{z}$ model.
-The equal-time spin correlations for the fitted $J_{123}$-$D_{z}$ model can be calculated using the self-consistent Gaussian approximation (SCGA) method Canals and Garanin (2001); sup .
-Since the critical correlations are underestimated in the SCGA method, a reduced $T=5$ K is assumed in the calculations to better describe the experimental data sup .
+At $T=6$ K, triangular shaped lobes are observed around the KK-($\frac{1}{3}$, $\frac{1}{3}$, 0) points.
+As indicated by the dashed curves, the shape of the spiral surface is reproduced by a $J_{1}^{h}$-$J_{2}^{h}$ honeycomb-lattice model with $J_{2}^{h}/J_{1}^{h}=0.72$, where $J_{1}^{h}$ $(J_{2}^{h})$ is equivalent to $J_{2}$ ($J_{3}$) in the $J_{1\textrm{-}5}$-$D_{z}$ model.
+The equal-time spin correlations for the fitted $J_{1\textrm{-}5}$-$D_{z}$ model can be calculated using the self-consistent Gaussian approximation (SCGA) method Canals and Garanin (2001); sup .
+Since the critical correlations are underestimated in the SCGA method, a reduced $T=5$ K is assumed in the calculations to better describe the experimental data.
 As compared in Fig.
 3(a), the calculated and experimental intensity distributions agree well with each other, both following the spiral surface of a SSL on a honeycomb lattice.
+Figure 3: (a) Left half shows the diffuse neutron scattering pattern in the ($h$, $k$, 0) plane for Cs3Fe2Cl9 measured at $T=6$ K on CORELLI.
+Data are symmetrized according to the $6/mmm$ Laue class.
+The right panel is the calculated diffuse neutron scattering pattern for the fitted $J_{1\textrm{-}5}$-$D_{z}$ model using the SCGA method assuming a reduced $T=5$ K to compensate for the underestimated critical correlations.
+Both the experimental and calculated data were integrated over l=[0.1,0.1]l=[-0.1,0.1] reciprocal lattice units, r.l.u.
+The solid-line hexagon indicates the boundary of the first Brillouin zone.
+Triangular-shaped lobes around the $K$-$(\frac{1}{3},\frac{1}{3},0)$ points are the spiral surface for a honeycomb-lattice model with a frustration ratio of $J_{2}^{\textrm{h}}/J_{1}^{\textrm{h}}=0.72$.
+(b) Similar experimental (left half) and calculated (right half) diffuse scattering patterns in the ($h$, $k$, 1) plane with an integration range of $l=[0.9,1.1]$ r.l.u.
+(c) Experimental (left half) and calculated (right half) diffuse scattering patterns with an integration range of l=[0.1,0.1]+[0.9,1.1]l=[-0.1,0.1]+[0.9,1.1] r.l.u.
+(d) Experimental (left half) and calculated (right half) diffuse scattering patterns in the $(h,h,l)$ plane with an integration width of 0.1 r.l.u.
+along (h,h,0)(h,-h,0).
+(e) Scattering intensity along $(h,h,0)$ integrated from the $(h,k,0)+(h,k,1)$ pattern in panel (c) with an integration width of 0.06 r.l.u.
+along (h,h,0)(h,-h,0) as outlined by the dashed-line rectangle in the left half of panel (c).
+(f) $l$-dependence of the scattering intensity integrated in the area of $[0.125,0.5]$ and [0.125,0.125][-0.125,0.125] r.l.u.
+along the $(h,h,0)$ and (h,h,0)(h,-h,0) directions, respectively.
+This area is outlined in the right half of panel (c) by a dashed-line rectangle.
+In panels (e) and (f), the red solid line shows the calculated results for the fitted $J_{1\textrm{-}5}$-$D_{z}$ model using the SCGA method.
 […]
-[…]
-[…]
+A special feature of the SSL in Cs3Fe2Cl9 is the visibility of the spiral surface, which arises from its unique codimension two.
+On the original honeycomb lattice with antiferromagnetic $J_{1}^{h}$, the spiral surface within the first Brillouin zone has a structure factor of zero due to the interference between the two sublattices Liu et al.
+(2016), leading to a diffuse pattern similar that shown in Fig.
+3(a).
+However, as compared in Figs.
+3(a) and (b), the visibility of the spiral surface in Cs3Fe2Cl9 is complementary between the $l=0$ and $l=1$ planes.
+This variance in visibility originates from the $l$-dependence of the phase factor, which modulates the interference between the two sublattices and thus reflects the higher codimension of the SSL Liu et al.
+(2016).
+By summing up the intensity in the planes of $l=0$ and 1, the complete spiral surface can be recovered in Fig.
+3(c), upon which a line cut along the (h,h,0h,~h,~0) direction is shown in Fig.
+3(e).
+The scattering pattern in the (h,h,lh,~h,~l) plane presented in Fig.
+3(d) reveals weak scattering intensity between the integer-$l$ planes.
+According to our calculations, the intensity of the interplanar scattering diminishes with decreasing temperatures, thus can be attributed to thermal excitations out of the ground state manifolds.
+This observation is further confirmed in Fig.
+3(f) through the measured and calculated $l$-dependence of the integrated intensity within a rectangular area of $[0.125,0.5]$ r.l.u.
+and [0.125,0.125][-0.125,0.125] r.l.u.
+along the $(h,h,0)$ and (h,h,0)(h,-h,0) directions, respectively.
 Competing orders and a possible ObD transition.
@@
 Two additional phases exist at $H\gtrsim 6$ T, including a $\frac{1}{2}$-magnetization plateau phase, VII, up to $\sim 12$ T and a transitional phase, VIII, that precedes the field-polarized FM phase at $H\gtrsim 15$ T.
-[…]
-Due to the limited number of magnetic reflections collected on WAND2, we resort to classical Monte Carlo simulations to understand the rich phases in Cs3Fe2Cl9.
-As explained in the Supplemental Material sup , the fitted $J_{123}$-$D_{z}$ model exhibits a relatively simple phase diagram that contains only two main phases with $\bm{q}_{\textrm{I}}=(\frac{1}{2},0,0)$.
-Therefore, starting from the fitted $J_{123}$-$D_{z}$ model, we examined the impacts of weak perturbations from isotropic exchange interactions up to the tenth neighbors one-by-one.
-The exchange paths of each perturbations and the representative theoretical phase diagrams are presented in the Supplemental Material sup .
-Following this method, we found that by adding a weak fifth-neighbor coupling $J_{5}=-0.008$ meV that only marginally impacts the INS spectra and diffuse scattering patterns sup , the calculated phase diagram, shown in Fig.
-4(f), reproduces the main phases observed in experiments.
-Phases VI and VI′′ (Phase V) adjacent to phases I and VII exhibit similar diffraction pattern as that in phase VI (V).
-These primed phases, V, VI, and VI′′, are absent in experiments and may arise from finite size effects in our simulations.
-The magnetic orders in the main phases are described in Figs.
-4(g)-(j), with their corresponding Ansatz and structure factor plots presented in the Supplemental Material sup .
-As is consistent with the experimental data, similar propagation vectors are observed in phases III-V.
-Among them, phases III and IV are determined to be of the spiral-type orders by inspecting the spin configuration snapshots in the Monte Carlo simulations, where the transition between phases III and IV can be attributed to the modulation of the $\bm{q}=0$ magnitude sup .
+Single crystal neutron diffraction experiments were performed on WAND2 Frontzek et al.
+(2018) at the High Flux Isotope Reactor (HFIR) at ORNL to clarify the LRO phases as a function of magnetic field up to 6 T sup .
+As summarized in Figs.
+4(e)-4(h), four different types of diffraction patterns are observed in the ($h$, $k$, 0) plane among phases I-VI.
+Compared to that in phase I, the diffraction pattern of phase II in Fig.
+4(f) exhibits additional weak magnetic Bragg peaks at $\bm{q}_{\textrm{II}}=(\frac{1}{4},\frac{1}{4},0)$, indicating the coexistence of minor $\uparrow\uparrow\uparrow\downarrow$ and major $\uparrow\downarrow\uparrow\downarrow$ magnetic domains, where $\uparrow$ ($\downarrow$) represents spins (antiparallel) parallel with $c$.
+Surprisingly, the diffraction patterns in phases III-V, among which the pattern for phase IV is shown in Fig.
+4(g), are similar to each other sup : all exhibiting main reflections at $\bm{q}_{\textrm{III}}=\frac{5}{6}\times(\frac{1}{3},\frac{1}{3},0)$ close to the center of the edges of the triangular-shaped spiral surface as indicated by dashed lines, together with weaker secondary reflections at $2\bm{q}_{\textrm{III}}$.
+The existence of the latter often suggests a field-distorted spiral or SDW phase Gignoux et al.
+(1998); Stüßer et al.
+(2002).
+As shown in Fig.
+4(h) and also compared in Fig.
+4(c) for the line cuts along the (h+1,h1,0)(h+1,h-1,0) direction, the $\bm{q}_{\textrm{III}}$ refelctions disappear in phase VI.
+Instead, new magnetic reflections belonging to $\bm{q}_{\textrm{VI}}=\frac{11}{9}\times(\frac{1}{3},\frac{1}{3},0)$ appear at the corners of the triangular-shaped spiral surface, which is exactly the position predicted by the ObD transition Mulder et al.
+(2010); Liu et al.
+(2020).
+Comparison of the line cuts along the (2h,4h,0)(-2h,4h,0) direction shown in Fig.
+4(d) reveal a clear shift in the $\bm{q}_{\textrm{VI}}$ and $2\bm{q}_{\textrm{III}}$ positions, evidencing the distinct magnetic order in phase VI.
+Using classical Monte Carlo simulations in a parallel tempering scheme as implemented in the SpinMC program Spi , we calculated the $H$-$T$ phase diagram for the fitted $J_{1\textrm{-}5}$-$D_{z}$ model.
+The calculated phase diagram, shown in Fig.
+4(b), agrees well with the experimental results shown in Fig.
+4(a), while some discrepancies existing in the transitional regime, including phases I’ and VI’ shadowed by dashed lines, may originate from the absence of quantum fluctuation in our calculations sup .
+The agreement between the experimental and calculated phase diagrams allows us to use the simulated magnetic structures as constraints in refining the magnetic orders against the neutron diffraction dataset collected in each phase sup .
+Following this method, the magnetic orders in phases I-VI can be determined, and the representative structures in the neighboring bilayers are shown in Figs.
+4(i)-(l).
+Phases III and IV are determined to be of an elongated spiral-type orders, and the transition between them can be attributed to the modulation of the $\bm{q}=0$ magnitude sup .
 As a contrast, both phases V and VI are revealed to be of the SDW orders with spins along the $c$ axis despite their different $\bm{q}$ vectors.
-In phases III-V where sufficient numbers of magnetic Bragg peaks can be extracted in our neutron diffraction experiments, the proposed magnetic structures agree with the experimental dataset sup , thus verifying the results of our Monte Carlo simulations.
 Since $\bm{q}_{\textrm{VI}}$ in phase VI is the predicted propagation vector of the ObD theory Mulder et al.
@@
 (2020), it is tantalizing to ascribe the transition from the SSL state to phase VI as entropy-driven.
-Compared to the theoretical phase diagram for the fitted $J_{123}$-$D_{z}$ model, it can be concluded that in the presence of a magnetic field, the $J_{5}$ perturbations favor orders with propagation vectors over the spiral surface, among which $\bm{q}_{\textrm{VI}}$ is stabilized by thermal fluctuations at higher temperatures.
 The proposed SDW character of phase VI, however, poses a challenge in accurately determining the free energy for states above $T_{N}$ under the current theoretical framework Bergman et al.
-(2007), making it difficult to exclude perturbations other than thermal fluctuations.
-Conclusion.
-Our neutron scattering experiments establish Cs3Fe2Cl9 as a host of codimension-two SSLs, where the high codimension is manifested through the phase tuning of the visibility of the spiral surface.
-By combining neutron diffraction and classical Monte Carlo simulations, we clarify the rich field-induced phases as competing spiral and SDW orders, revealing Cs3Fe2Cl9 as a candidate compound to study the ObD transition in frustrated magnets.
+(2007).
+Further studies will be required to clarify the character of the SSL-SDW phase transition.
+Discussion.
+From the perspective of experimental realization, the introduction of codimension into SSLs is not a mere classification for convenience but rather has important implications Yao et al.
+(2021).
+On the original honeycomb lattice, the realization of SSL requires a relatively strong second-neighbor interactions, which is rarely satisfied in real materials Gao et al.
+(2022a).
+However, as demonstrated in our work for Cs3Fe2Cl9, the relatively strong intralayer couplings on the AB-stacked triangular lattice, which correspond to the second-neighbor interactions on an effective honeycomb lattice, naturally favor the emergence of SSLs.
+Besides the many compounds that are isostructural with Cs3Fe2Cl9 Brüning et al.
+(2021); Leuenberger et al.
+(1986a, b, c), there are a rich variety of candidate systems like RRZnPO Wan et al.
+(2024); Nientiedt and Jeitschko (1998) and $R_{3}$Ru4Al12 Hirschberger et al.
+(2019); Gao et al.
+(2019); Kolincio et al.
+(2021) where the SSL physics might be achieved.
+Codimension-two SSLs also offer a plethora of novel physics that remain to be explored.
+One immediate question is the impacts of thermal and quantum fluctuations over the SSL state.
+Besides the thermally induced ObD transitions, there are theoretical discussions about the prospect of realization quantum spin liquid on the honeycomb lattice through SSLs Reuther et al.
+(2011); Niggemann et al.
+(2020).
+On the AB-stacked triangular lattice, this approach may still be valid, which resonates with the recent efforts of utilizing the further-neighbor interactions on a triangular lattice to achieve quantum spin liquids near the phase boundaries Gallegos et al.
+(2025).
+Another exciting direction to explore is the novel spin textures that may emerge from the codimension-two SSLs.
+On an AB-stacked triangular lattice, the inter- and intra-layer interactions may follow different symmetry constraints depending on the presence or location of the centrosymmetric points.
+This additional freedom in symmetry constraints can be highly beneficial for the realization of exotic spin textures like magnetic skyrmions Mohylna et al.
+(2022, 2025).
 Acknowledgements.
@@
 The vertical bars indicate the positions of the structural Bragg peaks for Cs3Fe2Cl9.
-The blue line at the bottom shows the difference of measured and calculated intensities.
+The blue line at the bottom shows the difference between measured and calculated intensities.
+Figure S2: Temperature evolution of the powder neutron diffraction data for Cs3Fe2Cl9 collected on POWGEN.
+Positions of the representative magnetic Bragg peaks $(\frac{1}{2}00)$, $(\frac{1}{2}01)$, and $(\frac{1}{2}02)$ are indicated.
 IV Neutron diffraction experiments on POWGEN
@@
 Refined crystal structure parameters are listed in Table S1.
-Due to the lacking of neutron diffraction data in the high-$Q$ region, we cannot reliable fit the thermal parameters.
-Therefore, a uniform thermal factor of $B_{\textrm{iso}}=0.1$2 is assumed for all atoms.
+Due to the lacking of neutron diffraction data in the high-$Q$ region, we cannot reliably fit the thermal parameters.
+Therefore, a uniform thermal factor of $B_{\textrm{iso}}=0.1$ Å2 is assumed for all atoms.
 Figure S2 plots the temperature evolution of the diffraction pattern.
@@
 This transition temperature is consistent with that observed in magnetic susceptibility.
-Figure S2: Temperature evolution of the powder neutron diffraction data for Cs3Fe2Cl9 collected on POWGEN.
-Positions of the representative magnetic Bragg peaks $(\frac{1}{2}00)$, $(\frac{1}{2}01)$, and $(\frac{1}{2}02)$ are indicated.
 The refinement result of the neutron diffraction pattern collected at $T=1.6$ K is shown in Fig.
@@
 Table S1: Refined crystal structure parameters for Cs3Fe2Cl9 at 15 K.
-The space group is $P6_{3}/mmc$ with lattice constants $a=7.1784(1)$ and $c=17.6374(3)$.
-A uniform thermal parameter of $B_{\textrm{iso}}=0.1$2 is assumed for all atoms.
+The space group is $P6_{3}/mmc$ with lattice constants $a=7.1784(1)$ and $c=17.6374(3)$ Å.
+A uniform thermal parameter of $B_{\textrm{iso}}=0.1$ Å2 is assumed for all atoms.
 The goodness-of-fit parameters are $R_{p}=18.6\%$ and $R_{wp}=10.2\%$.
@@
 Table S2: Refined crystal structure parameters for Cs3Fe2Cl9 at 1.6 K.
-The space group is $P6_{3}/mmc$ with lattice constants $a=7.1784(1)$ and $c=17.6369(2)$.
-A uniform thermal parameter of $B_{\textrm{iso}}=0.1$2 is assumed for all atoms.
+The space group is $P6_{3}/mmc$ with lattice constants $a=7.1784(1)$ and $c=17.6369(2)$ Å.
+A uniform thermal parameter of $B_{\textrm{iso}}=0.1$ Å2 is assumed for all atoms.
 The goodness-of-fit parameters are $R_{p}=15.2\%$ and $R_{wp}=11.3\%$.
@@
 V Inelastic neutron scattering experiments on CNCS
-Inelastic neutron scattering (INS) experiments on powder sample of Cs3Fe2Cl9 were performed on CNCS Ehlers et al.
+Inelastic neutron scattering (INS) experiments on powder samples of Cs3Fe2Cl9 were performed on CNCS Ehlers et al.
 (2011) at the SNS of the ORNL.
@@
 (2009).
-Figure S3: INS spectra for Cs3Fe2Cl9 measured on CNCS with an incident neutron energy of $E_{i}=1.0$ meV at $T=2$ K.
-Data were collected for approximately 8 hrs with the source operating at a power of 1.4 MW.
-Figure S3 presents the INS spectra for Cs3Fe2Cl9 measured with an incident neutron nergy of $E_{i}=1.0$ meV at $T=2$ K.
-An excitation gap of $\sim 0.2$ meV is resolved, which is consistent with the existence of uniaxial SIA in Cs3Fe2Cl9.
 VI Diffuse neutron scattering experiments on CORELLI
@@
 (2022).
-Figure S4 compares the diffuse scattering patterns calculated using the SCGA method for the fitted $J_{123}$-$D_{z}$ model at $T=5$ and 6 K.
-As explained in the main text, the SCGA method is a mean-field-based theory and may underestimate the critical scattering in systems with relatively strong SIA.
-If we assume $T=6$ K as that in experiments, the calculated scattering pattern will exhibit less contrast at momentum transfers over and out of the spiral surface.
-Therefore, a reduced temperature of $T=5$ K is assumed in Fig.
-3 of the main text to compensate for the underestimated critical scattering.
-On the other hand, the comparison between the calculations at $T=5$ and 6 K also reveals that the scattering intensity between the integer-$l$ planes is reduced at lower temperatures.
-This temperature dependence confirms thermal excitations as the origin of the interplanar scattering.
-Figure S4: Comparison between the diffuse scattering patterns calculated using the SCGA method for the fitted $J_{123}$-$D_{z}$ model at $T=5$ K (a-f) and 6 K (g-l).
-Panels (a-f) is a reproduction of Fig.
-3 in the main text.
-Figure S5: (a) Calculated diffuse scattering pattern in the ($h,~{}k,~{}0$) plane using the fitted $J_{123}$-$D_{z}$ model at $T=6$ K.
-(b) Temperature evolution of the calculated spin correlation function $\langle\bm{S}\cdot\bm{S}\rangle$ along ($h$, $h$, 0) around (0.5, 0.5, 0).
-The integration area is outlined in panel (a) by a dashed rectangle.
-Black dots indicate the peak positions at the corresponding temperatures through Gaussian fits.
-(c) Temperature evolution of the calculated spin correlations $\langle S_{x}\cdot S_{x}\rangle$ (red triangles), $\langle S_{z}\cdot S_{z}\rangle$ (yellow circles), and their ratio $\langle S_{x}\cdot S_{x}\rangle/\langle S_{z}\cdot S_{z}\rangle$ (black dots).
-The integration area is outlined in panel (a) by a dashed-line hexagon.
-Red line over the $\langle S_{x}\cdot S_{x}\rangle/\langle S_{z}\cdot S_{z}\rangle$ data is the fitting curve to the Arrhenius law $A_{1}\exp(-\frac{T-T_{N}}{\Delta_{\textrm{SIA}}})+A_{2}$ with fitted parameters of $A_{1}=-0.52(1)$, $\Delta_{\textrm{SIA}}=4.03$ K, and $A_{2}=0.90(1)$.
-VII Effects of the uniaxial anisotropy on the SSL
-Classical Monte Carlo simulations for the fitted $J_{123}$-$D_{z}$ model were performed to study the effects of the SIA on the SSL in Cs3Fe2Cl9.
-As a reference, the diffuse scattering pattern at $T=6$ K calculated by the SCGA method is shown in Fig.
-S5.
-Through classical Monte Carlo simulations that incorporate critical scattering near $T_{N}$, Fig.
-S5(b) summarizes the temperature evolution of the calculated scattering intensity along ($h+\frac{1}{2}$, $h+\frac{1}{2}$, 0), where the integration range is outlined by dashed rectangle in Fig.
-S5(b).
-For the fitted $J_{123}$-$D_{z}$ model, $T_{N}$ is found to be $\sim 5.5$ K.
-The separated peaks at $T/T_{N}>1$ correspond to thermally stabilized spiral correlations in the SSL state, while their gradually reduced separation with decreasing $T$ reveals the second-order character of the transition into the LRO.
-Figure S5(d) compares the in-plane ($x$) and out-of-plane ($z$) components of the calculated spin correlations as a function of $T$.
-Due to the existence of uniaxial SIA, the evolution of $\langle S_{x}\cdot S_{x}\rangle$ is nonmonotonic, and the ratio between the in-plane and out-of-plane components exhibits an Arrhenius-like behavior in a wide temperature regime.
-VIII Neutron diffraction experiments on WAND2
+VII Neutron diffraction experiments on WAND2
 Single crystal neutron diffraction measurements were performed using the WAND2 diffractometer Frontzek et al.
-(2018) at the High Flux Isotope Reactor HFIR with $\lambda=1.486$.
+(2018) at the High Flux Isotope Reactor HFIR with $\lambda=1.486$ Å.
 A crystal with mass of $\sim 7.5$ mg was aligned with the $c$ axis vertical, and then sealed in an aluminum can with helium exchange gas.
@@
 In phase VI, at $T=5.05$ K and $H=5.5$ T the sample was rotated over 90° in 0.1° steps with a total measurement time of 26 h.
-Figure S6: Single crystal neutron diffraction pattern for Cs3Fe2Cl9 in phases (a) III, (b) IV, and (c) V in the ($h$,$k$,0) plane.
+Figure S3: Single crystal neutron diffraction pattern for Cs3Fe2Cl9 in phases (a) III, (b) IV, and (c) V in the ($h$,kk,0) plane.
 The experimental conditions are listed in the text.
 The dashed rectangle emphasizes the area in which the relative intensity of the $\bm{q}_{\textrm{III}}$ and $2\bm{q}_{\textrm{III}}$ reflections are compared in the text.
+Figure S4: Comparison between the calculated and observed intensity of the nuclear reflections collected on WAND2.
 Figure S3 compares the neutron diffraction patterns in phases III, IV, and V.
@@
 As emphasized by the dashed rectangle in each panel, the intensity of the $2\bm{q}_{\textrm{III}}$ reflection in phase III is weaker than that in phase IV, which suggests higher field-induced magnetization in phase IV.
-Comparing phases IV and V, the $\bm{q}_{\textrm{III}}$ reflection becomes weaker in phase V while the $2\bm{q}_{\textrm{III}}$ reflection stays almost unchanged.
+As compared in Fig.
+4 of the main text for phases IV and V, the $\bm{q}_{\textrm{III}}$ reflection becomes weaker in phase V while the $2\bm{q}_{\textrm{III}}$ reflection stays almost unchanged.
 This evolution is consistent with a transition from a spiral structure in phase IV to a spin density wave (SDW) structure in phase V.
-IX Refinements of the nuclear and magnetic structures
+For a spiral order like that in phase IV, both the in-plane and out-of-plane components contribute to the intensity of the main $\bm{q}_{\textrm{III}}$ reflections, while only the out-of-plane component, under renormalization as described in the section of ‘Analytical expressions for the field-induced phases’, contributes to the $2\bm{q}_{\textrm{III}}$ reflections.
+For the collinear SDW order in phase V, the in-plane component becomes zero, thus the intensity of the main $\bm{q}_{\textrm{III}}$ reflections will be strongly reduced compared that in phase IV.
+VIII Refinements of the nuclear and magnetic structures
 Figure S4 compares the experimental intensity of the nuclear Bragg reflections observed on WAND2 to the calculations assuming the same crystal structure as that listed in Table S2.
@@
 The agreement between the observed and calculated intensity confirms the good quality of our crystal sample.
-Figure S7: Comparison between the calculated and observed intensity of the nuclear reflections collected on WAND2.
 Figure S5: Comparison between the calculated and observed intensity of the magnetic reflections collected on WAND2 for phase I (a), III (b), IV (c), and V (d).
-Figure S8 summarizes the fitting results of the magnetic reflections collected on WAND2 in phases I (a), III (b), VI (c), and V (d).
+Figure S5 summarizes the fitting results of the magnetic reflections collected on WAND2 in phases I (a), III (b), IV (c), and V (d).
 The measuring conditions are listed in the main text.
@@
 Considering the standard deviations, the fitted magnitude of the in-plane moment $M_{\parallel}$ is consistent with our classical Monte Carlo simulations, where phases III and IV exhibit an elongated helical structure, and phase V exhibits a spin density wave structure with only the $S_{z}$ components.
-X Classical Monte Carlo simulations
-Classical Monte Carlo simulations for the $J_{123}$-$D_{z}$ model
-$\mathcal{H}=\sum_{\langle ij\rangle\in n}J_{n}\bm{S}_{i}\cdot\bm{S}_{j}+D_{z}(S_{z})^{2}$ (S1)
-and its perturbed derivatives were performed using the SpinMC code that implements the single spin flip Metropolis algorithm bue .
-Unless alternately specified, a $12\times 12\times 4$ supercell with 2304 spins was employed in simulations.
-Observables including the heat capacity and magnetization were averaged over $10^{6}$ of measurement sweeps after $3\times 10^{5}$ thermalization sweeps, where each sweep represents 2304 attempted spin flips at randomly selected sites.
+IX Temperature evolution of the diffraction pattern
+Figure S6 presents the temperature evolution of the neutron diffraction pattern for Cs3Fe2Cl9 in the ($h$,kk,0) plane.
+Experiments were measured on the WAND2 ($T=1.5$ K) and CORELLI ($T=6$ and 10 K) diffractometers using the same piece of crystal.
+The enhanced diffuse scattering intensity at $T=6$ K confirms its magnetic origin, which is replaced by magnetic Bragg peaks at $T=1.5$ K.
+Figure S6: Temperature evolution of the neutron diffraction pattern in the ($h$,kk,0) plane.
+The pattern in panel (a) was measured on the WAND2 diffractometer.
+The patterns in (b) and (c) were measured on the CORELLI diffractometer, and data collected at $T=20$ K has been subtracted as background.
+X Modeling the spin dynamics
+To model the magnon dispersion observed in the CNCS experiments, we performed linear spin wave theory (LSWT) calculations using the SpinW software package Toth and Lake (2015).
+The powder-averaged dynamical structure factor, $S(Q,\omega)$, was calculated in steps of $0.05$ Å-1 over a momentum transfer range of $Q=0.35$ to $2.05$ Å-1 for the spectra collected with $E_{i}=3.32$ meV, and over the range of $Q=0.4$ to $0.95$ Å-1 for the spectra collected with $E_{i}=1.0$ meV.
+The calculated spectrum was convoluted with the theoretical instrumental energy resolution for each $E_{i}$ configuration plus an additional Gaussian broadening of 0.05 meV that accounts for thermal broadening.
+To reduce the impacts of neutron absorption, data below and above the energy transfer of $\Delta E=1.25$ meV in the $E_{i}=3.32$ meV spectra were treated separately with a fitted intensity ratio of about 3:2.
+The optimal exchange parameters were determined through a simultaneous fit for all three datasets including the low- and high-energy sections from the $E_{i}=3.32$ meV spectra and the $E_{i}=1$ meV spectra, using a weighting factor of 1:3:5, respectively.
+Figure S7: Fit to the INS data using the $J_{1\textrm{-}5}$-$D_{z}$ model with (a) $E_{\mathrm{i}}=$ 3 meV data and (b) $E_{\mathrm{i}}=$ 1 meV data.
+We began the INS fit with an initial $J_{123}$-$D_{z}$ model, which is the minimum model to realize spiral spin-liquids on the lattice of Cs3Fe2Cl9:
+$\mathcal{H}=\sum_{n=1,2,3}J_{n}\bm{S}_{i}\cdot\bm{S}_{j}+D_{z}(S_{z})^{2}\textrm{.}$ (S1)
+Starting from this minimum model that includes exchange interactions up to the third neighbors, we progressively includes further-neighbor exchange interactions to check whether they can further lower the $\chi^{2}$ factor.
+Through this procedure, we found that interactions up to the fourth neighbors, i.e., $J_{1}$ through $J_{4}$, are essential for accurately reproducing the INS spectra.
+The fifth-neighbor interactions, $J_{5}$, were also included in our model, as it slightly reduces the $\chi^{2}$ from 20.66 to 20.29, although its fitted strength of 0.0013(1) meV is already much weaker than the other terms.
+Further perturbations up to the seventh neighbors yield no improvement for the fit, and their fitted strengths, when included, are all in the orders of $10^{-4}$ meV or even less.
+We also investigated the impacts of symmetry-allowed anisotropic terms, including the Dzyaloshinskii-Moriya (DM) interactions or off-diagonal exchange terms, but no improvement was observed after introducing these terms in the fitting program.
+Therefore, in the final $J_{1\textrm{-}5}$-$D_{z}$ model, we considered only the isotropic exchange interactions up to the fifth neighbors.
+The fitting results for the INS spectra at representative $Q$ positions are summarized in Fig.
+S7.
+XI Classical Monte Carlo simulations
+Classical Monte Carlo simulations for the $J_{1\textrm{-}5}$-$D_{z}$ model were performed using the SpinMC code that implements the single spin flip Metropolis algorithm bue .
+Unless otherwise specified, a $16\times 16\times 4$ supercell with 4096 spins was employed in simulations.
+Observables, including the heat capacity and magnetization, were averaged over $10^{5}$ of measurement sweeps after $2\times 10^{5}$ thermalization sweeps, where each sweep represents 8192 attempted spin flips at randomly selected sites.
+An overrelaxation sweep was applied after each Monte Carlo sweep to reduce the autocorrelation.
 Magnetic structure factors were calculated through fast Fourier transform using the FFTW package fft .
-The parallel tempering algorithm was utilized to facilitate thermal equilibrium, which was performed simultaneously over 128 replicas on 128 cores with a geometric series of temperatures between 1 and 10 K.
-After every 10 Monte Carlo sweeps, a replica exchange was attempted and the successful exchange rates are above $\sim 25\%$ for all the neighboring replicas.
-X.1 The $J_{123}$ model
-Figure S9 summarizes the simulation results for the $J_{123}$ model with no SIA.
-The coupling strengths are $J_{1}=-0.238$, $J_{2}=0.085$, $J_{3}=0.053$ meV as fitted from the INS experimental data.
-In zero field, the $J_{123}$ model exhibits a long-range order transition at $T_{N}\sim 3.5$ K, which is lower than that observed in experiments.
-The phase diagram shown in Fig.
-S9(a) mainly consists of two phases.
-Magnetic structure factors shown in Figs.
-S9(b-d) with $\bm{q}\sim\frac{4}{3}\times(\frac{1}{3},\frac{1}{3},0)$ in low magnetic fields and $\bm{q}=(\frac{1}{2},\frac{1}{2},0)$ in high magnetic fields, which is very different from the experimental observations.
-Figure S9: (a) Theoretical phase diagram for the $J_{123}$ model calculated by the classical Monte Carlo simulations.
+The parallel tempering algorithm was utilized to facilitate thermal equilibrium, which was performed simultaneously over 80 replicas on 80 cores with a geometric series of temperatures between 1 and 10 K.
+After every 10 Monte Carlo sweeps, a replica exchange was attempted, and the successful exchange rates are above $\sim 25\%$ for all the neighboring replicas.
+XII Phase diagram for the $J_{1\textrm{-}5}$-$D_{z}$ model
+Figures S8-S10 summarize the simulation results for the $J_{1\textrm{-}5}$-$D_{z}$ model.
+By carefully analyzing the field and temperature dependence of the magnetic susceptibility (see Fig.
+S9) and specific heat (see Fig.
+S10), we obtain the phase diagram shown in Fig.
+S8(a).
+The typical magnetic structure factors shown in Figs.
+S8(b-d) reproduce the $\bm{q}_{\textrm{I}}=(\frac{1}{2},0,0)$ in phase I and VII at $H=$ 0 and 10 T, respectively, and $\bm{q}_{\textrm{III}}=\frac{5}{6}\times(\frac{1}{3},\frac{1}{3},0)$ in phase IV at $H=5$ T.
+Figure S8: (a) Theoretical phase diagram for the $J_{1\textrm{-}5}$-$D_{z}$ model calculated by the classical Monte Carlo simulations.
 Pseudocolor corresponds to the value of heat capacity in arbitrary units (arb.
 units).
-(b-d) Magnetic structure factors $\langle\bm{S}\cdot\bm{S}\rangle$ calculated at $T=1$ K in a field of (b) 0 T, (c) 5 T, and (d) 10 T.
-Calculations of the magnetic structure factor were performed on a $18\times 18\times 4$ supercell with $3\times 10^{5}$ thermalization sweeps and $5\times 10^{5}$ measurement sweeps.
-Figure S10: (a) Theoretical phase diagram for the $J_{123}+D_{z}$ model calculated by the classical Monte Carlo simulations.
-An easy-axis SIA of $D_{z}=-0.037$ meV is added to the $J_{123}$ model considered in Fig.
-S9.
-Pseudocolor corresponds to the value of heat capacity in arbitrary units (arb.
-units).
-(b-d) Magnetic structure factors $\langle\bm{S}\cdot\bm{S}\rangle$ calculated at $T=1$ K in a field of (b) 0 T, (c) 6 T, and (d) 10 T.
-Figure S11: Exchange paths for the perturbational couplings $J_{4}$ $J_{10}$ that are considered in our Monte Carlo simulations.
-X.2 The $J_{123}+D_{z}$ model
-Figure S10 summarizes the simulation results for the $J_{123}+{D_{z}}$ model that considers isotropic exchange interactions up to the third neighbors and a uniaxial SIA.
-The coupling strengths are $J_{1}=-0.238$, $J_{2}=0.085$, $J_{3}=0.053$, and $D_{z}=-0.037$ meV as fitted from the INS experimental data.
-In zero field, a magnetic long-range order transition is observed at $T_{N}\sim 5.8$ K, which is close to the experimental observation.
-Although the phase diagram shown in Fig.
-S10(a) mainly consists of two phases as that in Fig.
-S9(a), magnetic structure factors shown in Figs.
-S10(b-d) reveal that the propagation vector $\bm{q}$ stays at $(\frac{1}{2},\frac{1}{2},0)$ r.l.u.
-in both phases.
-This calculation indicates that the experimentally observed $\bm{q}_{\textrm{I}}=(\frac{1}{2},\frac{1}{2},0)$ r.l.u.
-in phase I is due to the existence of uniaxial SIA, yet the stabilization of the field-induced phases III, IV, V, and VI requires further perturbations in the spin Hamiltonian.
-Figure S12: (a) Theoretical phase diagram for the $J_{123}+D_{z}+J_{6}$ model calculated by the classical Monte Carlo simulations.
-A perturbative sixth-neighbor interaction of $J_{6}=0.01$ meV is added to the $J_{123}+D_{z}$ model considered in Fig.
-S10.
-Pseudocolor corresponds to the value of heat capacity in arbitrary units (arb.
-units).
-(b-d) Magnetic structure factors $\langle\bm{S}\cdot\bm{S}\rangle$ calculated at $T=1$ K in a field of (b) 0 T, (c) 5 T, and (d) 10 T.
-X.3 An overview of the perturbation effects
-As the theoretical phase diagram for the $J_{123}+D_{z}$ model cannot reproduce the experimental results, extensive classical Monte Carlo simulations were performed to explore possible perturbative terms that may stabilize the experimentally observed additional phases in intermediate fields.
-By confining the perturbations to a single-term Heisenberg exchange interaction up to the tenth neighbors, we calculated the phase diagram and magnetic structure factors for each perturbed model.
-The exchange paths for $J_{4}$-$J_{10}$ are indicated in Fig.
-S11.
-Figure S13: (a) Theoretical phase diagram for the $J_{123}+D_{z}+J_{5}$ model calculated by the classical Monte Carlo simulations.
-A perturbative fifth-neighbor interaction of $J_{5}=-0.008$ meV is added to the $J_{123}-D_{z}$ model considered in Fig.
-S10.
-Pseudocolor corresponds to the value of heat capacity in arbitrary units (arb.
-units).
-(b-d) Magnetic structure factors $\langle\bm{S}\cdot\bm{S}\rangle$ calculated at $T=1$ K in a field of (b) 0 T, (c) 5 T, and (d) 10 T.
-Calculations of the magnetic structure factor were performed on a $18\times 18\times 4$ supercell with $3\times 10^{5}$ thermalization sweeps and $5\times 10^{5}$ measurement sweeps.
+(b-d) Magnetic structure factors $\langle\bm{S}\cdot\bm{S}\rangle$ calculated at $T=1.5$ K in a field of (b) 0 T, (c) 5 T, and (d) 10 T.
+Calculations of the magnetic structure factor were performed on a $16\times 16\times 4$ supercell with $2\times 10^{5}$ thermalization sweeps and $1\times 10^{5}$ measurement sweeps.
 White arrows indicate the positions of weak magnetic Bragg peaks.
-Figure S14: Theoretical $M/H$ curves calculated by the classical Monte Carlo simulations for the $J_{123}+D_{z}+J_{5}$ model.
+Figure S11 compares the experimental and theoretical magnetization curves calculated by classical Monte Carlo simulations for the fitted $J_{1\textrm{-}5}$-$D_{z}$ model.
+The identified critical fields are largely consistent with those in the phase diagram in Fig.
+S8(a).
+The existence of the 1/2 magnetization plateau phase is well reproduced in our simulations.
+Figure S9: Theoretical magnetization curves calculated by the classical Monte Carlo simulations for the $J_{1\textrm{-}5}$-$D_{z}$ model.
 The strength of the applied field for each curve is indicated on the left.
@@
 (2021).
-Through our calculations, it was found that all of the considered perturbations with a strength lower than $\sim 0.01$ meV are able to introduce additional phases in the intermediate field regime, yet their effects on the magnetic structure factors categorize them into two groups.
-The first group contains antiferromagnetic (AFM) $J_{6}$ and $J_{7}$, or ferromagnetic (FM) $J_{8}$, $J_{9}$, and $J_{10}$ perturbations.
-However, in zero field, these types of perturbations do not produce the experimentally observed $\bm{q}_{\textrm{I}}=(\frac{1}{2},\frac{1}{2},0)$ in phase I and thus contradicts the experimental observations.
-The second group contains AFM $J_{4}$ or FM $J_{5}$ perturbations.
-This group of perturbations stabilize $\bm{q}_{\textrm{I}}=(\frac{1}{2},\frac{1}{2},0)$ in zero field while introducing additional phases similar to that observed in experiments.
-For both groups of perturbations, a sign reversal eliminates the field-induced phases, leading to phase diagrams similar to that of the unperturbed model shown in Fig.
-S10(a).
-Figure S15: Theoretical heat capacity $C_{p}$ calculated by the classical Monte Carlo simulations for the $J_{123}+D_{z}+J_{5}$ model in the field regime of (a) [0, 3], (b) [4, 7], and (c) [8, 14] T.
+Figure S10: Theoretical heat capacity $C_{p}$ calculated by the classical Monte Carlo simulations for the $J_{1\textrm{-}5}$-$D_{z}$ model in the field regime of (a) [0, 3.2], (b) [4, 7], and (c) [8, 14] T.
 Data are normalized by the maximal $C_{p}$ value, $C_{p}^{\textrm{max}}$.
@@
 (2021).
-X.4 The $J_{123}+D_{z}+J_{6}$ model
-As an illustration for the first group of perturbations, Fig.
-S10 summarizes the simulation results for the $J_{123}+{D_{z}}+J_{6}$ model that considers a sixth-neighbor AFM perturbation of $J_{6}=0.01$ meV.
-Comparison of the phase diagrams shown in Fig.
-S10(a) and Fig.
-S12(a) reveals that the perturbation of $J_{6}$ causes additional phases in the intermediate field regime between $\sim 4.5$ and 6.5 T.
-However, as shown in Fig.
-S12(b-d), the propagation vector of the long-range ordered phase in zero field deviates from $\bm{q}_{\textrm{I}}=(\frac{1}{2},\frac{1}{2},0)$ r.l.u.
-and thus fails to reproduce the experimental results.
-Similar effects are observed for the remaining perturbations in the first group, thus excluding them as the main perturbation to the $J_{123}+{D_{z}}$ Hamiltonian.
-X.5 The $J_{123}+D_{z}+J_{5}$ model
-As discussed in the previous section, AFM $J_{4}$ and FM $J_{5}$ belong to the second group of perturbations and have similar effects on the phase diagram.
-However, there is one slight difference for these two perturbations: In zero field, AFM $J_{4}$ and FM $J_{5}$ raises and lowers $T_{N}$, respectively.
-This observation favors FM $J_{5}$ since the experimental $T_{N}\sim 5.4$ K is lower than the theoretically predicted $T_{N}\sim 5.8$ K for the unperturbed $J_{123}+D_{z}$ model.
-Figures S13-S15 summarize the simulation results for the $J_{123}+{D_{z}}+J_{5}$ model that considers a FM perturbation of $J_{5}=-0.008$ meV.
-This strength of $J_{5}$ is found to best describe the experimental phase diagram on a $12\times 12\times 4$ supercell.
-By carefully analyzing the field and temperature dependence of the magnetic susceptibility (see Fig.
-S14) and specific heat (see Fig.
-S15), we obtain the phase diagram shown in Fig.
-S13(a).
-The magnetic structure factors shown in Figs.
-S13(b-d) reproduce the $\bm{q}_{\textrm{I}}=(\frac{1}{2},\frac{1}{2},0)$ in zero field and $\bm{q}_{\textrm{III}}=\frac{5}{6}\times(\frac{1}{3},\frac{1}{3},0)$ at $H=5$ T.
-Therefore, we conclude that FM $J_{5}$ is the main perturbative term to the $J_{123}+{D_{z}}$ Hamiltonian, leading to a $J_{123}+{D_{z}}+J_{5}$ model that is employed in the main text.
-Figure S16: Theoretical magnetic structure factors (a) $\langle S_{z}S_{z}\rangle$ and (b) $\langle S_{x}S_{x}\rangle$ in phase III for the $J_{123}+D_{z}+J_{5}$ model.
-Calculations were performed on a $15\times 15\times 4$ supercell using the classical Monte Carlo simulations.
-The strength of $J_{5}$ is slightly increased to $-0.01$ meV to compensate for the size effect so that the phase boundaries are similar to that shown in Fig.
-4(f) of the main text.
-The temperature and magnetic field are set to 1 K and 4 T, respectively.
-Calculations were performed over $5\times 10^{5}$ measuring sweeps after $5\times 10^{5}$ thermalization sweeps.
+Figure S11: Experimental (a) and theoretical (b) magnetization curves.
+The experimental data, together with the phase boundaries indicated by the dashed lines, are adapted from Ref.
+Ishii et al.
+(2021).
+The theoretical data were obtained from classical Monte Carlo simulations using the $J_{1\textrm{-}5}$-$D_{z}$ model.
+Figure S12: Theoretical magnetic structure factors (a) $\langle S_{z}S_{z}\rangle$ and (b) $\langle S_{x}S_{x}\rangle$ in phase III for the $J_{1\textrm{-}5}$-$D_{z}$ model.
+Calculations were performed on a $16\times 16\times 4$ supercell using the classical Monte Carlo simulations.
+The temperature and magnetic field are set to 2.33 K and 4 T, respectively.
+Calculations were performed over $1\times 10^{5}$ measuring sweeps after $2\times 10^{5}$ thermalization sweeps.
 (c-h) Similar magnetic structure factors in phases IV, V, and VI.
-The calculation parameters are $T=1$ K and $H=5.5$ T in phase IV, $T=3.1$ K and $H=4.0$ T in phase V, and $T=5.0$ K and $H=4.0$ T in phase VI.
-X.5.1 Spiral-type orders in phases III and IV
+The calculation parameters are $T=2.33$ K and $H=5$ T in phase IV, $T=4.17$ K and $H=4.0$ T in phase V, and $T=5.58$ K and $H=5.0$ T in phase VI.
+XII.1 Spiral-type orders in phases III and IV
 Comparisons of the direction-resolved magnetic structure factors and real-space spin configurations unveil the magnetic structures in each of the field-induced phases.
@@
 In phases III and IV, both the $\langle S_{z}S_{z}\rangle$ and $\langle S_{x}S_{x}\rangle$ components exhibit long-range order, suggesting a spiral-type magnetic order that involves both in-plane and out-of-plane spin components.
-The magnetic structure factor along the $z$ direction is about 10 times higher than that along the $x$ direction, which indicates that the spiral order in phases III and IV is elongated along the $c$ axis due to the uniaxial SIA.
-Compared to that in phase III, the relatively stronger $2\bm{q}_{\textrm{III}}$ reflections in phase IV suggests stronger squaring-up effects in a higher magnetic field.
+The structure factor along the $z$ direction is about 10 times higher than that along the $x$ direction, which indicates that the spiral order in phases III and IV is elongated along the $c$ axis due to the uniaxial SIA.
+Compared to that in phase III, the relatively stronger $2\bm{q}_{\textrm{III}}$ reflections in phase IV suggest stronger squaring-up effects in a higher magnetic field.
 This field dependence of the $2\bm{q}_{\textrm{III}}$ reflections is consistent with the experimental observations shown in Fig.
-S6.
-X.5.2 SDW orders in phases V and VI
+S3.
+XII.2 SDW orders in phases V and VI
 As shown in Fig.
-S16(f) and (h), the $\langle S_{x}S_{x}\rangle$ component of the magnetic structure factors in phases V and VI exhibit diffuse patterns, while sharp Bragg peaks are observed in the $\langle S_{z}S_{z}\rangle$ component.
+S12(f) and (h), the $\langle S_{x}S_{x}\rangle$ component of the magnetic structure factors in phases V and VI exhibits diffuse patterns, while sharp Bragg peaks are observed in the $\langle S_{z}S_{z}\rangle$ component.
 This observation indicates that magnetic moments in phases V and VI are ordered only along the $c$ axis, leading to sinusoidally modulated SDW orders in these two phases.
 We also note that comparison of the $\langle S_{z}S_{z}\rangle$ component in phases III (Fig.
-S16(a)), IV (Fig.
-S16(c)), and V (Fig.
-S16(e)) reveals a similar magnetic propagation vector $\bm{q}_{\textrm{III}}$ and its high harmonics $2\bm{q}_{\textrm{III}}$, which reproduces the experimental observations shown in Fig.
-S16.
-XI Analytical expressions for the field-induced phases
+S12(a)), IV (Fig.
+S12(c)), and V (Fig.
+S12(e)) reveals a similar magnetic propagation vector $\bm{q}_{\textrm{III}}$ and its high harmonics $2\bm{q}_{\textrm{III}}$, which reproduces the experimental observations shown in Fig.
+S3.
+XII.3 Magnetic orders in phases VII and VIII
+A few possible magnetic orders have been proposed for the 1/2-magnetization plateau phase Ishii et al.
+(2021).
+Through classical Monte Carlo simulations, the order in the 1/2-magnetization plateau phase of the $J_{1\textrm{-}5}$-$D_{z}$ model is presented in Fig.
+S13(a).
+In the same figure, the predicted magnetic order in phase VIII of the $J_{1\textrm{-}5}$-$D_{z}$ model is presented in Fig.
+S13(b).
+Further neutron diffraction experiments in fields above 6 T will be required to verify these magnetic orders.
+Figure S13: Magnetic orders in phases (a) VII, (b) VIII, and (c) I’ viewed along the $c$ axis as determined from classical Monte Carlo simulations for the $J_{1\textrm{-}5}$-$D_{z}$ model.
+The bottom left (top right) part depicts the spin configuration for two sublattices (one sublattice).
+XII.4 Additional intermediate phases
+As compared in Fig.
+4 of the main text, the calculated phase diagram contains two intermediate phases that are possibly not observed in the experiment.
+In the regime of the experimental phase II, the calculated structure is a double-$\mathbf{q}$ phase that is denoted as I’, where the ordering vectors are the same as those in phase I.
+The simulated magnetic order in this phase is presented in Fig.
+S13(c).
+Close to the lower phase boundary of the 1/2-plateau phase VII, another transitional phase, designated VI’, is identified in the calculated phase diagram.
+The magnetic order in phase VI’ is largely the same as that of phase VI, while the $\langle S_{x}S_{x}\rangle$ and $\langle S_{y}S_{y}\rangle$ components is enhanced in phase VI’ due to its relatively low temperatures.
+In our classical Monte Carlo simulations, we did not manage to remove these additional phases, especially phase VI’, by introducing further perturbations to the $J_{1\textrm{-}5}$-$D_{z}$ model.
+Therefore, we propose that the absence of this transitional phase in experiments might be due to quantum fluctuation that is not considered in our classical Monte Carlo calculations.
+On the antiferromagnetic triangular lattice model, quantum fluctuation is known to stabilizes a 1/3 magnetization plateau with a collinear up-up-down magnetic order even in the case of $S=5/2$ Coletta et al.
+(2016).
+Similarly, in Cs3Fe2Cl9, quantum fluctuation may enhance the stability of the 1/2 magnetization plateau phase and thus remove the transitional VI’ phase obtained in classical Monte Carlo simulations.
+XIII Analytical expressions for the field-induced phases
 To further verify the magnetic structures of the field-induced phases, we calculate the magnetic structure factors for the following ansatz on a honeycomb lattice to describe the spiral-type orders in phases III and IV (Eqn.
@@
 (S3)):
-$\bm{M}(\bm{r})=\left\{\begin{array}[]{l}M_{\perp}\cos(\bm{q}\cdot\bm{r})\bm{n}_{1}+M_{\parallel}\sin(\bm{q}\cdot\bm{r})\bm{n}_{2}+M_{z}\bm{n}_{1}\textrm{,\qquad\qquad\ \ if}\ \bm{r}\in\{r_{1}\}\\ M_{\perp}\cos(\bm{q}\cdot\bm{r}+\phi)\bm{n}_{1}+M_{\parallel}\sin(\bm{q}\cdot\bm{r}+\phi)\bm{n}_{2}+M_{z}\bm{n}_{1}\textrm{,\quad if}\ \bm{r}\in\{r_{2}\}\end{array}\right.$ (S2)
-$\bm{M}(\bm{r})=\left\{\begin{array}[]{l}M_{\perp}\cos(\bm{q}\cdot\bm{r})\bm{n}_{1}+M_{z}\bm{n}_{1}\textrm{,\qquad\ \ if}\ \bm{r}\in\{r_{1}\}\\ M_{\perp}\cos(\bm{q}\cdot\bm{r}+\phi)\bm{n}_{1}+M_{z}\bm{n}_{1}\textrm{,\quad if}\ \bm{r}\in\{r_{2}\}\end{array}\right.$ (S3)
+$\bm{M}(\bm{r})=\left\{\begin{array}[]{l}M_{\perp}\cos(\bm{q}\cdot\bm{r})\bm{n}_{1}+M_{\parallel}\sin(\bm{q}\cdot\bm{r})\bm{n}_{2}+M_{z}\bm{n}_{1}\textrm{,\qquad\qquad\ \ if}\ \bm{r}\in\{r_{1}\}\\ M_{\perp}\cos(\bm{q}\cdot\bm{r}+\phi)\bm{n}_{1}+M_{\parallel}\sin(\bm{q}\cdot\bm{r}+\phi)\bm{n}_{2}+M_{z}\bm{n}_{1}\textrm{,\quad if}\ \bm{r}\in\{r_{2}\}\end{array}\right$.
+(S2)
+$\bm{M}(\bm{r})=\left\{\begin{array}[]{l}M_{\perp}\cos(\bm{q}\cdot\bm{r})\bm{n}_{1}+M_{z}\bm{n}_{1}\textrm{,\qquad\ \ if}\ \bm{r}\in\{r_{1}\}\\ M_{\perp}\cos(\bm{q}\cdot\bm{r}+\phi)\bm{n}_{1}+M_{z}\bm{n}_{1}\textrm{,\quad if}\ \bm{r}\in\{r_{2}\}\end{array}\right$.
+(S3)
 Figure S14: (a) Theoretical magnetic structure factors (a) $\langle S_{z}S_{z}\rangle$ and (b) $\langle S_{x}S_{x}\rangle$ using the ansatz in Eqn.
 (S2) for phase III.
-Calculations were performed for a two-dimensional $30\times 30$ honeycomb superlattice through Fourier transform.
+Calculations were performed for a two-dimensional $30\times 30$ honeycomb superlattice through the Fourier transform.
 (c-d) Magnetic structure factors using the ansatz in Eqn.
@@
 By comparing the magnetic structure factors in Fig.
-S17 with those in Fig.
-S16, Fig.
-S6, and Fig.
+S14 with those in Fig.
+S12, Fig.
+S3, and Fig.
 4 in the main text, it is confirmed that the double peaks along over the Brillouin zone boundaries are due to high harmonics $2\bm{q}_{\textrm{III}}$ reflections induced by magnetic field, thus corroborating the proposed spiral-type and SDW orders in the field-induced phases.
-XII Magnetic order in the magnetization plateau phase
-A few possible magnetic orders have been proposed for the 1/2-magnetization plateau phase Ishii et al.
-(2021).
-Through classical Monte Carlo simulations, the order in the 1/2-magnetization plateau phase of the $J_{123}+{D_{z}}+J_{5}$ model is presented in Fig.
-S18.
-Further neutron diffraction experiments in fields above 6 T will be required to verify this magnetic order.
-Figure S18: Magnetic order in phase VII viewed along the $c$ axis as determined from classical Monte Carlo simulations for the $J_{123}+{D_{z}}+J_{5}$ model.
-The bottom left (top right) part depicts the spin configuration for two sublattices (one sublattice).
-XIII Impacts of the $J_{5}$ perturbation on the spin dynamics
-Figure S19 presents the calculated INS spectra and diffuse neutron scattering pattern in the $(h,k,0)$ plane for the $J_{123}+D_{z}+J_{5}$ model with a perturbative $J_{5}=-0.008$ meV.
-The INS spectra are calculated by the linear spin wave theory.
-The diffuse neutron scattering pattern is calculated by the SCGA method at $T=5$ K.
-The calculated results only exhibit marginal difference compared to those of the $J_{123}+D_{z}$ model as presented in the main text.
-Figure S19: (a) Calculated INS spectra and (b) diffuse neutron scattering pattern in the $(h,k,0)$ plane for the $J_{123}+D_{z}+J_{5}$ model with $J_{5}=-0.008$ meV.
-The diffuse neutron scattering pattern is calculated by the SCGA method at $T=5$ K.
+Figure S15: (a) Calculated diffuse scattering pattern in the (h,k,0h,~k,~0) plane using the fitted $J_{123}$-$D_{z}$ model at $T=6$ K.
+(b) Temperature evolution of the calculated spin correlation function $\langle\bm{S}\cdot\bm{S}\rangle$ along ($h$, $h$, 0) around (0.5, 0.5, 0).
+The integration area is outlined in panel (a) by a dashed rectangle.
+Black dots indicate the peak positions at the corresponding temperatures through Gaussian fits.
+(c) Temperature evolution of the calculated spin correlations $\langle S_{x}\cdot S_{x}\rangle$ (red triangles), $\langle S_{z}\cdot S_{z}\rangle$ (yellow circles), and their ratio $\langle S_{x}\cdot S_{x}\rangle/\langle S_{z}\cdot S_{z}\rangle$ (black dots).
+The integration area is outlined in panel (a) by a dashed-line hexagon.
+Red line over the $\langle S_{x}\cdot S_{x}\rangle/\langle S_{z}\cdot S_{z}\rangle$ data is the fitting curve to the Arrhenius law $A_{1}\exp(-\frac{T-T_{N}}{\Delta_{\textrm{SIA}}})+A_{2}$ with fitted parameters of $A_{1}=-0.561(4)$, $\Delta_{\textrm{SIA}}=3.93(7)$ K, and $A_{2}=0.913(2)$.
+XIV Effects of the uniaxial anisotropy on the SSL
+Classical Monte Carlo simulations for the fitted $J_{1\textrm{-}5}$-$D_{z}$ model were performed to study the effects of the SIA on the SSL in Cs3Fe2Cl9.
+As a reference, the diffuse scattering pattern at $T=6$ K calculated by the SCGA method is shown in Fig.
+S15.
+Through classical Monte Carlo simulations that incorporate critical scattering near $T_{N}$, Fig.
+S15(b) summarizes the temperature evolution of the calculated scattering intensity along ($h+\frac{1}{2}$, $h+\frac{1}{2}$, 0), where the integration range is outlined by dashed rectangle in Fig.
+S15(b).
+For the fitted $J_{1\textrm{-}5}$-$D_{z}$ model, $T_{N}$ is found to be $\sim 5.75$ K.
+The separated peaks at $T/T_{N}>1$ correspond to thermally stabilized spiral correlations in the SSL state, while their gradually reduced separation with decreasing $T$ reveals the second-order character of the transition into the LRO.
+Figure S15(c) compares the in-plane ($x$) and out-of-plane ($z$) components of the calculated spin correlations as a function of $T$.
+Due to the existence of uniaxial SIA, the evolution of $\langle S_{x}\cdot S_{x}\rangle$ is nonmonotonic, and the ratio between the in-plane and out-of-plane components exhibits an Arrhenius-like behavior with an activation gap of $\Delta_{\textrm{SIA}}=3.93(7)$ K in a wide temperature regime.
+The value of the activation energy is close to the estimation from the single ion anisotropic energy of $D_{z}S(S+1)\sim 3.27$ K that is required to flip a single spin from the easy axis along $c$ to the $ab$ plane.